### v1

# Magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances in hot and dense strange hadronic matter

Abstract
This work comprises the calculations of magnetic moments of $\frac{1}{2}^{-}$ baryon resonances in the presence of hot and dense hadronic matter. Using the chiral $SU(3)$ quark mean field model, we have taken into account the impact on in-medium scalar meson fields to investigate the density effects on the in-medium baryon masses and their constituent quarks. In light of chiral constituent quark model $\chi$CQM, we have calculated the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances and scrutinized the explicit contributions coming from the valence quarks, sea quarks and the orbital angular momentum of sea quarks. The effective baryonic magnetic moments have been investigated for the varying values of isospin asymmetry and strangeness fraction.
I Introduction
In quantum chromodynamics (QCD), quarks and gluons are treated as fundamental degrees of freedom which form bound states known as hadrons (baryons($qqq$) and mesons($q\bar{q}$)). These fundamental degrees of freedom undergo dynamical changes due to strong interactions in dense nuclear medium, which in turn cause in-medium modifications in the properties of hadrons. Confinement makes it challenging to understand QCD at lower energy scales. Scrutinizing the internal structure and the properties of light hadrons allow us to unravel the physics of strong interactions at low energies. Apart from heavy-ion collision experimental studies on chiral symmetry restoration at high temperature thirtyseven , and the de-confinement of hadrons to quark-gluon plasma (QGP) thirtyeight , it has become a primary interest for experiments such as PANDA at FAIR thirtythree ; thirtyfour , BESIII at BEPCII thirtysix ; thirtytwo , HADES at GSI thirtyfive etc., to explore the electromagnetic properties of hadrons in dense matter. Electromagnetic form factors, being key observable to investigate the electric and magnetic moment distributions, are of theoretical interest to explore other primary hadronic properties like magnetic moment and charged radii to probe deep inside of hadrons.
Magnetic moment of hadrons, in free space, play a very important role in investigating their internal structure at a sub-nuclear level sixtyfour . Theoretically, the magnetic moments of octet as well as decuplet baryons have been extensively studied in free space one ; two ; three ; four ; five ; six ; seven ; eight ; nine . Isgur framed his model with one-gluon exchange, in order to account for internal structure of hadrons isgur1 ; isgur2 ; isgur3 ; isgur4 . The constituent quark model (CQM) propounded that the baryonic magnetic moments can be calculated as the summation of the individual magnetic moments of its constituent quarks ten ; eleven . It was Glozman who incorporated the Goldstone boson exchange (GBE) interaction in the low energy regime and the concept of one-gluon exchange was substituted by vector meson exchange interactions, which in fact led to better fits for baryon spectra glozman . It was not until the development of chiral constituent quark model ($\chi\mathrm{CQM}$) that the pseudo-scalar and scalar meson exchanges were considered which was later extended to the inclusion of vector meson exchange dai . Hence, $\chi\mathrm{CQM}$ was formed by extending the quark model and further incorporating the chiral symmetry and corresponding spontaneous symmetry breaking into it twelve . The $\chi\mathrm{CQM}$ model provides satisfactory explanation to the “Proton spin puzzle”, which was an outcome of the final data taking at EMC, and other hadron properties ten ; thirteen ; fourteen ; fifteen ; sixteen ; seventeen ; eighteen ; nineteen ; twenty ; twentyone ; twentytwo ; twentythree ; twentyfour ; twentyfive ; twentysix ; twentyseven . In previous studies, this model has been used for calculating the baryonic magnetic moments, by taking into account the sea quark polarizations and their orbital angular momentum contribution through the Cheng-Li mechanism twentyeight ; twentynine ; thirty . In the $\chi$CQM formalism, the magnetic moments of low-lying ${\frac{1}{2}}^{-}$ $N^{*}$ resonances thirtyone , ${\frac{1}{2}}^{-}$ octet baryon resonances thirtyone and low-lying ${\frac{1}{2}}^{-}$ $\Lambda$ resonances A_Torres have been calculated in earlier studies. Apart from $\chi$CQM, the magnetic moments of $N^{*}$ resonances have been computed by three quark Hamiltonian approach and the results were not consistent with lattice QCD values with very large variations hamiltonian_approach . As an extension to chiral perturbation theory, the Chiral unitary model was also used in ref. cum_hyodo to calculate the magnetic moments of $N^{*}(1535)$ resonances.
As mentioned earlier, the study of magnetic moments of baryons and the medium dependent effects on the electromagnetic properties is an area of active research and stands relevant to some of the experiments. In order to study the properties of nuclear matter, various theoretical models like the MIT bag model thirtynine , linear sigma model Petro , non-linear sigma model fortyfive , Zimanyi and Moszkowski model seventyone , cloudy bag model fortythree , Nambu-Jona-Lasinio (NJL) model fortytwo , light-front quark model Schlumpf1993 , parity doublet model paritydoublet1 ; paritydoublet2 etc., have been developed. The cloudy bag model attained increased accuracy over MIT bag model in explaining QCD effects by incorporating the presence of pion clouds at the core of a baryon. With further development, models like the NJL model successfully explained low energy baryonic dynamics by considering chiral quarks as constituents of hadrons, through the inclusion of the interaction between the constituent quarks njlsuccess . Few years later, the quark meson coupling (QMC) model was constituted and it facilitated a deeper understanding of the in-medium baryon properties fortyfour . The QMC model, formulated in lines of cloudy bag and NJL model, has previously been deployed to study the impact of medium on the magnetic moments of octet baryons and has produced results which are in good agreement with experimental data seventyfour . The chiral $SU(3)$ quark mean field model (CQMF) fortyfour ; fortyfive ; fortysix ; fortyseven is an extended version of the quark mean field model fortyeight , which is based on the QMC model approach. In the CQMF model, the constituent quarks are confined into baryons by a confining potential based on mean field approximation using classical expectation values instead of quantum field operators fortynine . In the context of medium modifications of electromagnetic properties of hadrons, the studies in ref. kaon show the in-medium effects on $K^{+}$ meson properties in light-front quark meson coupling model (LF-QMC) in symmetric nuclear matter. With the rise in medium density, the kaon electromagnetic form factors decrease faster than in vacuum, while the kaon electromagnetic charge radii grow in size. By studying pion and kaon meson structures in NJL model the authors in ref. njlmeson have obtained results consistent with that in ref. kaon . For octet baryons in light of the QMC model effoctet , it has been shown that the in-medium electric form factors are dampened and the magnetic form factors are enhanced, and this finding is consistent with the in-medium variations of the bound proton electromagnetic form factors observed at Jefferson Lab jlab . The study of effective magnetic moments of baryons was also extended to calculate the low-lying charm, and the low-lying bottom baryons with the aim of understanding the roles of the heavy and light quarks in symmetric nuclear matter tsushima1 ; tsushima2 . While a comparative study on octet baryons with modified quark meson coupling model (MQMC) mqmc , suggests that the observed in-medium baryonic magnetic moments incur drastic variations compared to that in QMC framework. In CQMF formalism, the medium modifications of magnetic moments of octet fiftythree , decuplet fiftyfour and charmed Dutt:2024lui baryons have been studied in hot and dense hadronic matter. The CQMF model has recently been used to study valence quark distribution and transverse momentum dependent distributions of valence quarks of pions Puhan:2024xdq ; Kaur:2024wze .
Presence of strangeness fraction ($f_{s}$) is a principal aspect to understand the departure of hadronic properties of baryons in strange hadronic matter comprising of hyperons from that in a non-strange medium. The in-medium variations of magnetic moments of octet and decuplet baryons in strange medium has shown stark deviation, as studied in fiftythree ; fiftyfour ; sixtytwo , for baryons of higher strangeness quantum number. In the present work, we have focused our study on the in-medium effects on the magnetic moments of the low-lying $N^{*}$ resonances and $\frac{1}{2}^{-}$ octet baryon resonances in CQMF model. It is exciting to study these resonances, as their detection connects hadron models to QCD properties in low and intermediate energy regimes. Photoproduction has been a key experimental tool involved in detecting the low-lying states of baryon resonances as shown in Refs. fifty ; fiftyone ; fiftytwo . The large decay widths of these excited baryon resonances, makes it challenging to detect them in experiments. In our current work, we have incorporated $\chi\mathrm{CQM}$ to calculate the properties of the baryon resonances as in thirtyone and have visualized them as a system of three valence quarks surrounded by the meson cloud. We have calculated the in-medium magnetic moments of the ${\frac{1}{2}}^{-}$ baryon resonances at finite temperature and density of hadronic medium through the medium modification of the corresponding constituent quark masses and the baryon masses. Within the CQMF framework, the effective masses of quarks ($m^{*}_{q}$) are calculated through the medium modified values of the scalar fields $\sigma$ and $\zeta$ and scalar iso-vector field $\delta$. The confined quarks inside the baryons, participate in the exchange interaction with these scalar meson fields as well as vector meson fields $\omega$, $\rho$ and $\phi$.
This manuscript has been organized as follows: In section II the CQMF model used to compute the in-medium quark masses and effective baryon masses, at finite temperature and density of the medium is presented. In section III we have calculated the contributions of valence quarks, sea quarks and orbital angular momentum of sea quarks towards the total magnetic moments of baryons, in light of $\chi\mathrm{CQM}$ model. Section IV is dedicated to a detailed discussion on the results obtained from numerical calculations. We devote section V to present the summary of the current work.
II Chiral Quark Mean Field Model
In this section, we discuss briefly the framework, CQMF model, to calculate the in-medium constituent quark and baryon masses. As discussed earlier, in the CQMF model the quarks and mesons are treated as fundamental degrees of freedom, and the quarks are assumed to be confined inside baryons by an effective confining potential. The constituent quarks and mesons obtain their masses through the inclusion of properties such as spontaneous and explicit (for pseudoscalar mesons) symmetry breaking fortyfour . In the CQMF model, the effective Lagrangian density accounting for various interaction terms is given as fortyfour ; fortyfive ; fiftythree
${\cal L}_{{\rm eff}}\,=\,{\cal L}_{q0}\,+\,{\cal L}_{qm}\,+\,{\cal L}_{VV}\,+\,{\cal L}_{\Sigma\Sigma}\,+\,{\cal L}_{\chi SB}\,+\,{\cal L}_{c},$ (1)
where $\,{\cal L}_{q0}=\bar{q}i\gamma^{\mu}\partial_{\mu}q\,$ is the kinetic term for the free part of massless quarks and ${\cal L}_{qm}\,$ is the quark-meson interaction term that remains invariant under chiral $SU(3)$ transformations. ${\cal L}_{qm}\,$ can be written as fortyfour ; fortysix ; fiftythree ; fiftyfour
$\mathcal{L}_{qm}$ $=g_{\mathrm{s}}\left(\bar{\Psi}_{L}M\Psi_{R}+\bar{\Psi}_{R}M^{\dagger}\Psi_{L}\right)-g_{v}\left(\bar{\Psi}_{L}\gamma^{\mu}l_{\mu}\Psi_{L}+\bar{\Psi}_{R}\gamma^{\mu}r_{\mu}\Psi_{R}\right)$ (2)
$=\frac{g_{\mathrm{s}}}{\sqrt{2}}\bar{\Psi}\left(\sum_{a=0}^{8}s_{a}\lambda_{a}+\mathrm{i}\gamma^{5}\sum_{a=0}^{8}p_{a}\lambda_{a}\right)\Psi-\frac{g_{v}}{2\sqrt{2}}\bar{\Psi}\left(\gamma^{\mu}\sum_{a=0}^{8}v_{\mu}^{a}\lambda_{a}-\gamma^{\mu}\gamma^{5}\sum_{a=0}^{8}a_{\mu}^{a}\lambda_{a}\right)\Psi.$
In eq. 2, $\Psi=$ $\begin{pmatrix}u\\ d\\ s\end{pmatrix},$ and $g_{\mathrm{s}}$ and $g_{v}$ are the respective scalar and vector coupling constants. The relations between various quark-meson coupling constants are given as fortysix
$\frac{g_{s}}{\sqrt{2}}=$ $g_{\delta}^{u}=-g_{\delta}^{d}=g_{\sigma}^{u}=g_{\sigma}^{d}=\frac{1}{\sqrt{2}}g_{\zeta}^{s},~{}~{}~{}~{}~{}g_{\delta}^{s}=g_{\sigma}^{s}=g_{\zeta}^{u}=g_{\zeta}^{d}=0\,,$ (3)
$\frac{g_{v}}{2\sqrt{2}}=g_{\rho}^{u}=-g_{\rho}^{d}=g_{\omega}^{u}=g_{\omega}^{d},~{}~{}~{}~{}~{}~{}~{}g_{\omega}^{s}=g_{\rho}^{s}=0.$ (4)
It is important to note that spin-0 scalar ($\Sigma$) and pseudoscalar ($\Pi$) meson nonets can be expressed as fortyfour ; fortysix
$M\left(M^{\dagger}\right)=\Sigma\pm i\Pi=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}\left(s^{a}\pm i\pi^{a}\right)\lambda^{a},$ (5)
where $s^{a}$ and $\pi^{a}$ represent the nonets of scalar and pseudoscalar mesons, respectively, while $\lambda^{a}$ are Gell-Mann matrices with $\lambda^{0}=\sqrt{\frac{2}{3}}I$. The spin-1 mesons can also be defined in a similar way as
$l_{\mu}\left(r_{\mu}\right)=\frac{1}{2}\left(V_{\mu}\pm A_{\mu}\right)=\frac{1}{2\sqrt{2}}\sum_{a=0}^{8}\left(v_{\mu}^{a}\pm a_{\mu}^{a}\right)\lambda^{a}.$ (6)
In the above expression, $v_{\mu}^{a}$ and $a_{\mu}^{a}$ are nonets of vector and pseudovector mesons, respectively. The physical states of scalar and vector meson nonets are given by
$\Sigma=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}s_{a}\lambda^{a}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\sigma+\delta^{0}\right)&\delta^{+}&\kappa^{*+}\\ \delta^{-}&\frac{1}{\sqrt{2}}\left(\sigma-\delta^{0}\right)&\kappa^{*0}\\ \kappa^{*-}&\bar{\kappa}^{*0}&\zeta\end{array}\right),$ (10)
and
$V_{\mu}=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}v_{\mu}^{a}\lambda^{a}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\omega_{\mu}+\rho_{\mu}^{0}\right)&\rho_{\mu}^{+}&K_{\mu}^{*+}\\ \rho_{\mu}^{-}&\frac{1}{\sqrt{2}}\left(\omega_{\mu}-\rho_{\mu}^{0}\right)&K_{\mu}^{*0}\\ K_{\mu}^{*-}&\bar{K}_{\mu}^{*0}&\phi_{\mu}\end{array}\right),$ (14)
respectively.
The term, ${\cal L}_{VV}$ in eq. 1 represents the chiral-invariant vector meson self interaction terms. Within the framework of mean field approximation ${\cal L}_{VV}$ is given by fortyfour ; fiftythree ; fiftyfour
${\cal L}_{VV}=\frac{1}{2}\,\frac{\chi^{2}}{\chi_{0}^{2}}\left(m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2}+m_{\phi}^{2}\phi^{2}\right)+g_{4}\left(\omega^{4}+6\omega^{2}\rho^{2}+\rho^{4}+2\phi^{4}\right).$ (15)
For the scalar mesons, their self interaction terms are written as (fourth term in eq. 1)
${\cal L}_{\Sigma\Sigma}=$ $-\frac{1}{2}\,k_{0}\chi^{2}\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)+k_{1}\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)^{2}+k_{2}\left(\frac{\sigma^{4}}{2}+\frac{\delta^{4}}{2}+3\sigma^{2}\delta^{2}+\zeta^{4}\right)$
$+k_{3}\chi\left(\sigma^{2}-\delta^{2}\right)\zeta-k_{4}\chi^{4}-\frac{1}{4}\chi^{4}{\rm ln}\frac{\chi^{4}}{\chi_{0}^{4}}+\frac{\xi}{3}\chi^{4}{\rm ln}\left(\left(\frac{\left(\sigma^{2}-\delta^{2}\right)\zeta}{\sigma_{0}^{2}\zeta_{0}}\right)\left(\frac{\chi^{3}}{\chi_{0}^{3}}\right)\right).$ (16)
The parameters $k_{0},k_{1},k_{2},k_{3}$ and $k_{4}$ appearing in the above Lagrangian density are calculated using masses of $\pi$ and $K$ as well as the average mass of $\eta$ and $\eta^{{}^{\prime}}$ mesons. The vacuum expectation values of scalar meson fields $\sigma$ and $\zeta$, i.e. $\sigma_{0}$ and $\zeta_{0}$, can be expressed in terms of pion and kaon leptonic decay constants $f_{\pi}$ and $f_{K}$, through the relations
$\sigma_{0}=-f_{\pi}~{}~{}{\rm and}~{}~{}~{}~{}\zeta_{0}=\frac{1}{\sqrt{2}}\left(f_{\pi}-2f_{K}\right),$ (17)
where $f_{\pi}=92.8$ MeV and $f_{K}=115$ MeV. The vacuum value of dilaton field $\chi_{0}=254.6$ MeV and the coupling constant $g_{4}=37.4$, obtained from fit reasonable effective nucleon mass fortyfour ; fiftythree ; fiftyfour . The last three terms of eq. 16 are introduced in the model to incorporate the trace anomaly property of QCD, which leads to the trace of energy momentum tensor being proportional to fourth power of dilaton field $\chi$ fortynine . The parameter $\xi$, originating from the logarithmic term used in eq. 16, is usually determined from QCD $\beta$-function at one loop level for three colors and flavors fortynine . The fifth term of the effective Lagrangian density of eq. 1 is the explicit chiral symmetry breaking term, ${\cal L}_{\chi SB}$, and is expressed as
${\cal L}_{\chi SB}=\frac{\chi^{2}}{\chi_{0}^{2}}\left[m_{\pi}^{2}f_{\pi}\sigma+\left(\sqrt{2}\,m_{K}^{2}f_{K}-\frac{m_{\pi}^{2}}{\sqrt{2}}f_{\pi}\right)\zeta\right].$ (18)
The above expression is introduced in chiral effective models to account for the non-vanishing pseudoscalar meson masses and it satisfies the partial conserved axial-vector current relations for $\pi$ and $K$ mesons fortyfour ; fortysix ; fortyseven . The confinement of constituent quarks inside the baryons, due to the presence of a confining scalar-vector potential $\chi_{c}$, is introduced through the last term in the Lagrangian density, in eq. 1, and is written as
${\cal L}_{c}=-\bar{\psi}\chi_{c}\psi,$ (19)
where the confining potential $\chi_{c}$ is given by
$\chi_{c}(r)=\frac{1}{4}k_{c}\,r^{2}(1+\gamma^{0})\,.$ (20)
In eq. 20 above the coupling constant $k_{c}$ is taken to be $98\,\text{MeV}.\text{fm}^{-2}$.
In order to investigate the properties of asymmetric strange matter at finite temperature and density, we are using mean field approximation fortyfour . Under the influence of meson mean field, we express the Dirac equation for the quark field $\Psi_{qi}$ as
$\left[-i\vec{\alpha}\cdot\vec{\nabla}+\chi_{c}(r)+\beta m_{q}^{*}\right]\Psi_{qi}=e_{q}^{*}\Psi_{qi}.$ (21)
Here the quark $q$ ($q=u,d,s$) is in a baryon of type $i$ ($i=N,\Lambda,\Sigma,\Xi$)  and $\vec{\alpha}$ , $\beta$  are the usual Dirac matrices. The effective quark mass $m_{q}^{*}$ is defined as
$m_{q}^{*}=-g_{\sigma}^{q}\sigma-g_{\zeta}^{q}\zeta-g_{\delta}^{q}I^{3q}\delta+m_{q0},$ (22)
At zero baryon density, the masses of constituent quarks $u$, $d$, $s$ are expressed in terms of vacuum expectation values $\sigma_{0}$ and $\zeta_{0}$ of meson scalar fields $\sigma$ and $\zeta$, respectively. For the light $u$ and $d$ quarks, we have
$m_{u}=m_{d}=-g_{\sigma}^{q}\sigma_{0}=-\frac{g_{s}}{\sqrt{2}}\sigma_{0}.$ (23)
The values of coupling constant $g_{s}$ and confining parameter $k_{c}$ are fitted to obtain binding energy of $-16.0$ MeV at nuclear saturation density, $\rho_{0}=0.16$ $\text{fm}^{-3}$. This fit provides the value of effective mass of light quarks to be equal to $256$ MeV. Here $m_{q0}$ is zero for non-strange $u$ and $d$ quarks, whereas for strange $s$ quark $m_{q0}=\Delta m=77$ MeV. The term $\Delta m$ is introduced through the Lagrangian density ${\cal L}_{\Delta m}=-\Delta m\bar{\psi}m_{1}\psi$, where $S_{1}=\frac{1}{3}\left(1-\lambda_{8}\sqrt{3}\right)$. In eq. 21, $e_{q}^{*}$ in the Dirac equation is the effective energy of quarks and is expressed as fortyfour
$e_{q}^{*}=e_{q}-g_{\omega}^{q}\omega-g_{\rho}^{q}I^{3q}\rho-g_{\phi}^{q}\phi.$ (24)
The in-medium mass of a baryon is expressed in terms of its effective energy $E^{*}_{i}$ and spurious center of mass momentum $p_{icm}$ as in Barik1985 ; Barik2013 ,
$M^{*}_{i}=\sqrt{E^{*2}_{i}-<p^{*2}_{icm}>}.$ (25)
Further the in-medium energies of $i^{th}$ baryon can be related to the in-medium energies of quarks and number of q type constituent quarks $n_{qi}$ in $i^{th}$ baryon by the expression
$E^{*}_{i}=\sum_{q}n_{qi}e_{q}^{*}+E_{ispin}.$ (26)
Here, $E_{ispin}$ is a correction term in the effective energy of a baryon that accounts for spin-spin interaction between constituent quarks. The $E_{ispin}$ is determined so as to obtain the correct baryon masses at zero density. Finally, the thermodynamic potential $\Omega$ describing the isospin asymmetric strange hadronic matter at finite temperature and baryon density is written as fortyfour
$\Omega=-\frac{k_{B}T}{(2\pi)^{3}}\sum_{i}\gamma_{i}\int_{0}^{\infty}d^{3}k\left\{\ln\left(1+e^{-\left[E_{i}^{*}(k)-\nu_{i}^{*}\right]/k_{B}T}\right)+\ln\left(1+e^{-\left[E_{i}^{*}(k)+\nu_{i}^{*}\right]/k_{B}T}\right)\right\}-\mathcal{L}_{M}-\mathcal{V}_{\text{vac }},$ (27)
where the mesonic Lagrangian is $\mathcal{L}_{M}=\mathcal{L}_{VV}+\mathcal{L}_{\Sigma\Sigma}+\mathcal{L}_{\chi SB}$ and we have $E^{*}_{i}(k)=\sqrt{M_{i}^{*2}+k^{2}}$. In the above equation, the degeneracy factor is $\gamma_{i}=2$ for baryons and the summation is over hyperons and nucleons. In eq. 27, the effective chemical potential $\nu_{i}^{*}$ is expressed in terms of the usual baryon chemical potential $\mu_{i}$ as fortyfour ; fiftythree ; fiftyfour ; sixtytwo
$\nu_{i}^{*}=\mu_{i}-g_{\phi}^{i}\phi-g_{\omega}^{i}\omega-g_{\rho}^{i}I^{3i}\rho.$ (28)
From eq. 27 the equations of motion of scalar and vector meson fields are calculated by minimizing the effective thermodynamic potential as (see APPENDIX)
$\frac{\partial\Omega}{\partial\sigma}=\frac{\partial\Omega}{\partial\zeta}=\frac{\partial\Omega}{\partial\chi}=\frac{\partial\Omega}{\partial\delta}=\frac{\partial\Omega}{\partial\omega}=\frac{\partial\Omega}{\partial\rho}=\frac{\partial\Omega}{\partial\phi}=0.$ (29)
Equation 29 is solved for finite values of temperature and density of the medium. In order to account for isospin asymmetry in the hadronic medium a parameter $i_{a}=-\frac{\Sigma_{i}I_{3i}\rho_{i}}{\rho_{B}}$ is introduced, where $I_{3i}$ represents the third component of isopin and $\rho_{B}$ is baryonic density of the medium fiftyfour . To take into account hyperons in the medium, we define strangeness fraction $f_{s}=\frac{\Sigma_{i}\left|s_{i}\right|\rho_{i}}{\rho_{B}}$, here $|s_{i}|$ denotes the number of $s$ quarks in $i^{th}$ baryon.
We have eq. 25 which addresses our aim to compute the effective baryon masses, by taking into account the effective constituent quark masses, baryon density and the spurious center of mass motion. In the upcoming section we shift the trajectory of our discussion towards the calculation of effective baryonic magnetic moments, by incorporating the medium modified constituent quark masses and the effective baryon masses.
III Magnetic Moment of ${\frac{1}{2}}^{-}$ baryon resonances
In the present section, methodology used to calculate the magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances has been discussed. The formalism of chiral constituent quark model ($\chi\mathrm{CQM}$) thirtyone , has been used to calculate the magnetic moment of the baryons through the following fluctuation process eight ; fiftyfive
$q^{\uparrow\downarrow}\rightarrow\mathrm{GB}+q^{\prime\downarrow\uparrow}\rightarrow{\left(q\overline{q}^{\prime}\right)+q^{\prime\downarrow\uparrow}},$ (30)
where GB represents the emitted Goldstone boson and $q\overline{q}^{\prime}+q^{\prime}$ constitute the sea quarks nineteen ; twenty ; twentyone ; twentytwo ; twentythree . The $\chi\mathrm{CQM}$ model incorporates the concept of confinement and chiral symmetry breaking twelve ; fiftysix , where the massless quarks acquire mass through spontaneous breaking of chiral symmetry twelve . In this approach we consider the emission of a GB, which further splits into a $q\overline{q}^{\prime}$ pair. The coupling between the quark and GBs is described by the effective interaction Lagrangian,
$\mathcal{L}=g_{8}\overline{\mathbf{q}}\left(\Phi+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}I\right)\mathbf{q}=g_{8}\overline{\mathbf{q}}\left(\Phi^{\prime}\right)\mathbf{q}.$ (31)
This quark-GB coupling represented in this Lagrangian is weak enough to consider the fluctuation in eq. 30 as a small perturbation and the higher order fluctuations can therefore be neglected. The process of emission of GB from the quark gives rise to the probability of a spin-flip of the quark. In eq. 31, $\zeta^{\prime}=g_{1}/g_{8},$ where $g_{1}$ and $g_{8}$ denote the coupling constants for the singlet and octet GBs, respectively and $I$ is a $3\times 3$ identity matrix. The matrix of the GBs can be written as
$\Phi^{\prime}=\left(\begin{array}[]{ccc}\frac{\pi^{0}}{\sqrt{2}}+\beta\frac{\eta}{\sqrt{6}}+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}&\pi^{+}&\alpha K^{+}\\ \pi^{-}&-\frac{\pi^{0}}{\sqrt{2}}+\beta\frac{\eta}{\sqrt{6}}+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}&\alpha K^{0}\\ \alpha K^{-}&\alpha\bar{K}^{0}&-\beta\frac{2\eta}{\sqrt{6}}+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}\end{array}\right)$ (32)
and
$q=\left(\begin{array}[]{l}u\\ d\\ s\end{array}\right).$
The $SU(3)$ chiral symmetry breaking is introduced in $\chi\mathrm{CQM}$ by considering that $m_{s}>m_{u,d}$ and the masses of GBs as non-degenerate $\left(m_{\eta^{\prime}}>m_{K,\eta}>m_{\pi}\right)$ fiftyseven . The parameter $a=\left|g_{8}\right|^{2}$ is denoted as the transition probability of chiral fluctuation $u(d)\rightarrow$ $d(u)+\pi^{+(-)}$, whereas $a\alpha^{2},a\beta^{2}$ and $a\zeta^{\prime 2}$ denote the probabilities of transitions of $u(d)\rightarrow s+K^{-(0)},u(d,s)\rightarrow$ $u(d,s)+\eta$ and $u(d,s)\rightarrow u(d,s)+\eta^{\prime}$, respectively nineteen ; twenty ; twentyone ; twentytwo ; twentythree . By thoroughly analysing the spin and flavor distribution functions of proton, the best fitted set of parameters crucial to our calculations are nineteen ; twentythree ; twentyfive ; twentyeight ; twentynine ; thirty ; thirtyone : $a=0.12,\alpha=\beta=0.45,\eta=-0.15$.
The effective baryonic magnetic moment $\mu^{*}_{B}$ is a summation of the contributions due to spin polarization $\mu^{*S}_{B}$ and the angular momentum polarization $\mu^{*L}_{B}$, expressed as fiftythree ; fiftyfour
$\mu^{*}_{B}=\mu^{*S}_{B}+\mu^{*L}_{B}.$ (33)
Supposedly B ¿ is a SU(6) spin-flavor wave function of a baryon, then we have
$\left\langle B\left|\mu^{*S}_{B}\right|B\right\rangle=\Delta u\mu^{*}_{u}+\Delta d\mu^{*}_{d}+\Delta s\mu^{*}_{s},$ (34)
$\left\langle B\left|\mu^{*L}_{B}\right|B\right\rangle=\Delta u^{(1)}\mu^{*}_{u}+\Delta d^{(1)}\mu^{*}_{d}+\Delta s^{(1)}\mu^{*}_{s}.$ (35)
Here, $\mu^{*}_{q}(q=u,d,s)$ is the effective quark magnetic moment. The spin and orbital angular momentum polarizations are quantitatively comprehended from $\Delta q=q^{\uparrow}-q^{\downarrow}$ and $\Delta q^{(1)}=q^{(1)}-q^{(-1)}$, respectively. The polarizations can be calculated using the number operator $\mathcal{N}$ (=$\mathcal{N}^{\uparrow\downarrow}+\mathcal{N}^{\pm 1}$) as
$\widehat{B}\equiv\langle B|\mathcal{N}|B\rangle.$ (36)
Here $\mathcal{N}^{\uparrow\downarrow}$ and $\mathcal{N}^{(\pm 1)}$ of the number operator $\mathcal{N}$ are expressed as
$\mathcal{N}^{\uparrow\downarrow}=$ $n_{u\uparrow}u^{\uparrow}+n_{u\downarrow}u^{\downarrow}+n_{d\uparrow}d^{\uparrow}+n_{d^{\downarrow}}d^{\downarrow}+n_{s^{\uparrow}}s^{\uparrow}+n_{s\downarrow}s^{\downarrow},$ (37)
$\mathcal{N}^{(\pm 1)}=$ $n_{u^{(1)}}u^{(1)}+n_{u(-1)}u^{(-1)}+n_{d^{(1)}}d^{(1)}+n_{d^{(-1)}}d^{(-1)}+n_{s^{(1)}}s^{(1)}+n_{s^{(-1)}}s^{(-1)},$ (38)
with $n_{q^{\uparrow}}\left(n_{q^{\downarrow}}\right)$ denoting the number of spin up (down) quarks and $n_{q^{(1)}}\left(n_{q^{(-1)}}\right)$ being the number of quarks with the orbital angular momentum projection $m_{L}=1$ $\left(m_{L}=-1\right)$.
Following the Cheng Li mechanism ten , the calculation of total baryonic magnetic moment involves the contribution of valence and sea quark spin polarizations and orbital angular momentum of sea quarks. The spin part in eq. 33, can be obtained from contributions due to valence, sea and orbital angular momentum of sea quarks and is given by
$\mu^{*S}_{B}=\mu^{*S}_{val}+\mu^{*S}_{sea}+\mu^{*S}_{orbit}.$ (39)
The $\mu^{*S}_{val}$, of a baryon, can be derived by calculating the magnetic moments for the physical eigenstates of the orbital angular momentum $L=1$ ($P-$wave) negative-parity resonances thirtyone . For each baryon, the contribution of sea quark spin polarization in spin part of the total magnetic moment is obtained by substituting for each valence quark
$q^{\uparrow(\downarrow)}\rightarrow-P_{[q,GB]q^{\uparrow(\downarrow)}}+\left|\psi\left(q^{\uparrow(\downarrow)}\right)\right|^{2}.$ (40)
In eq. 40, $P_{[q,GB]}$ is the probability of emission of GBs from a quark $q^{\uparrow(\downarrow)}$ fiftyeight given by
$P(u\rightarrow\mathrm{GB})=P(d\rightarrow\mathrm{GB})=\frac{a}{6}\left(9+6\alpha^{2}+\beta^{2}+2\zeta^{\prime 2}\right),$ (41)
$P(s\rightarrow\mathrm{GB})=\frac{a}{3}\left(6\alpha^{2}+2\beta^{2}+\zeta^{\prime 2}\right),$
and $\left|\psi\left(q^{\uparrow(\downarrow)}\right)\right|^{2}$ represents the probability of transforming a $q^{\uparrow(\downarrow)}$ quark and it is given by nineteen ; twenty ; twentyone ; twentytwo ; twentythree ; fiftythree ; fiftyfour
$\left|\psi\left(u^{\uparrow(\downarrow)}\right)\right|^{2}=\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)u^{\uparrow(\downarrow)}+ad^{\uparrow(\downarrow)}+a\alpha^{2}s^{\uparrow(\downarrow)},$ (42)
$\left|\psi\left(d^{\uparrow(\downarrow)}\right)\right|^{2}=au^{\uparrow(\downarrow)}+\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)d^{\uparrow(\downarrow)}+a\alpha^{2}s^{\uparrow(\downarrow)},$
$\left|\psi\left(s^{\uparrow(\downarrow)}\right)\right|^{2}=a\alpha^{2}u^{\uparrow(\downarrow)}+a\alpha^{2}d^{\uparrow(\downarrow)}+\frac{a}{3}\left(2\beta^{2}+\zeta^{\prime 2}\right)s^{\uparrow(\downarrow)}.$
The orbital angular momentum contribution due to sea quarks, is expressed as ten ; fourteen ; twentyeight ; twentynine ; thirty ; thirtyone
$\mu^{*}\left(q^{\uparrow}\rightarrow q^{\prime\downarrow}\right)=\frac{e_{q}}{2m^{*}_{q}}\left\langle l_{q}\right\rangle+\frac{e_{q}-e_{q\prime}}{2M_{\mathrm{GB}}}\left\langle l_{\mathrm{GB}}\right\rangle,$ (43)
where $l_{q}\right>=\frac{M_{\mathrm{GB}}}{M_{\mathrm{GB}}+m^{*}_{q}}$ and $l_{\mathrm{GB}}\right>=\frac{m^{*}_{q}}{M_{\mathrm{GB}}+m^{*}_{q}}$. The quantities $\left(l_{q},l_{\mathrm{GB}}\right)$ and $\left(m^{*}_{q},M_{\mathrm{GB}}\right)$ stands for the orbital angular momenta and masses of quarks and GBs, respectively. The orbital moment of each process in eq. 43, when multiplied by the probability of occurrence of the process, produces the magnetic moment due to all possible transitions for a given valence quark. This can be mathematically expressed as,
${\left[\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\left(\frac{1}{2}+\frac{\beta^{2}}{6}+\frac{\zeta^{\prime 2}}{3}\right)\mu^{*}\left(u_{\uparrow}\rightarrow u_{\downarrow}\right)\right.$ (44)
$\left.+\mu^{*}\left(u_{\uparrow}\rightarrow d_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(u_{\uparrow}\rightarrow s_{\downarrow}\right)\right],$
${\left[\mu^{*}\left(d_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\mu^{*}\left(d_{\uparrow}\rightarrow u_{\downarrow}\right)+\left(\frac{1}{2}+\frac{\beta^{2}}{6}+\frac{\zeta^{\prime 2}}{3}\right)\right.$
×$\left.\times\mu^{*}\left(d_{\uparrow}\rightarrow d_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(d_{\uparrow}\rightarrow s_{\downarrow}\right)\right],$
${\left[\mu^{*}\left(s_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\alpha^{2}\mu^{*}\left(s_{\uparrow}\rightarrow u_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(s_{\uparrow}\rightarrow d_{\downarrow}\right)\right.$
$\left.+\left(\frac{2}{3}\beta^{2}+\frac{\zeta^{\prime 2}}{3}\right)\mu^{*}\left(s_{\uparrow}\rightarrow s_{\downarrow}\right)\right].$
The orbital moments of the $u,d$ and $s$ quarks can also be expressed in terms of the $\chi\mathrm{CQM}$ parameters $(a,\alpha,\beta,\zeta^{\prime})$, masses of GBs $\left(M_{\pi},M_{\eta\prime},M_{K},M_{\eta}\right)$ and effective quark masses $\left(m^{*}_{u},m^{*}_{d},m^{*}_{s}\right)$, respectively as thirtyone ; fiftythree ; fiftyfour :
${\left[\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\frac{3m_{u}^{*2}}{2M_{\pi}\left(m_{u}^{*}+M_{\pi}\right)}-\frac{\alpha^{2}\left(M_{K}^{2}-3m_{u}^{*2}\right)}{2M_{K}\left(m_{u}^{*}+M_{K}\right)}\right.+\frac{\beta^{2}M_{\eta}}{6\left(m_{u}^{*}+M_{\eta}\right)}$
$\left.+\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{3\left(m_{u}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{u},$
${\left[\mu^{*}\left(d_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{3\left(M_{\pi}^{2}-2m_{d}^{*2}\right)}{4M_{\pi}\left(m^{*}_{d}+M_{\pi}\right)}-\frac{\alpha^{2}M_{K}}{2\left(m_{d}^{*}+M_{K}\right)}-\frac{\beta^{2}M_{\eta}}{12\left(m_{d}^{*}+M_{\eta}\right)}\right.$ (45)
$\left.-\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{6\left(m_{d}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{d},$
${\left[\mu^{*}\left(s_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{\alpha^{2}\left(M_{K}^{2}-3m_{s}^{*2}\right)}{2M_{K}\left(m_{s}^{*}+M_{K}\right)}-\frac{\beta^{2}M_{\eta}}{3\left(m_{s}^{*}+M_{\eta}\right)}-\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{6\left(m_{s}^{*}+M_{\eta^{\prime}}\right)}\right]\mu_{s}^{*}.$
In the present study, we have included the effect of relativistic quark confinement and have calculated the mass adjusted effective magnetic moments of its constituent quarks by the expression fiftyfive
$\mu_{d}^{*}=-\left(1-\frac{\Delta M}{M_{B}^{*}}\right),\quad\mu_{s}^{*}=-\frac{m_{u}^{*}}{m_{s}^{*}}\left(1-\frac{\Delta M}{M_{B}^{*}}\right),\quad\mu_{u}^{*}=-2\mu_{d}^{*}.$ (46)
In the above relations, $\Delta M=M_{B,vac}-M_{B}^{*}$ where $M_{B,vac}$ is the vacuum mass of the given baryon and $M_{B}^{*}$ denotes the in-medium mass of this baryon.
In light of this formalism, we can obtain the spin contribution ($\mu_{B}^{*S}$) towards the baryonic magnetic moment ($\mu^{*}_{B}$). The expressions for spin contribution of n* resonance, of $\frac{1}{2}^{-}$ baryon octet, is obtained as in thirtyone
$\mu_{\text{val}}^{*S}\left(n^{*}\right)=-\left(\frac{2}{9}\mu_{d}^{*}+\frac{1}{9}\mu_{u}^{*}\right)\cos^{2}\theta+\left(\frac{10}{9}\mu_{d}^{*}+\frac{5}{9}\mu_{u}^{*}\right)\sin^{2}\theta\\ -\frac{8}{9}\left(\mu_{d}^{*}-\mu_{u}^{*}\right)\sin\theta\cos\theta,$
$\hskip 15.6491pt\mu_{\text{sea }}^{*S}\left(n^{*}\right)=$ $\frac{a}{9}\left[\left(5+2\alpha^{2}+\frac{2\beta^{2}}{3}+\frac{4\zeta^{\prime 2}}{3}\right)\mu_{d}^{*}\right.\left.+\left(4+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\mu_{u}^{*}+3\alpha^{2}\mu_{s}^{*}\right]\cos^{2}\theta$
$-\frac{5a}{9}\left[\left(5+2\alpha^{2}+\frac{2\beta^{2}}{3}+\frac{4\zeta^{\prime 2}}{3}\right)\right.\left.\mu_{d}^{*}+\left(4+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\mu_{u}^{*}+3\alpha^{2}\mu_{s}^{*}\right]\sin^{2}\theta$
$+\frac{8a}{9}\left(1+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\left(\mu_{d}^{*}-\mu_{u}^{*}\right)\sin\theta\cos\theta,$
$\mu_{\text{orbit }}^{*S}\left(n^{*}\right)=$ $-\left(\frac{2}{9}\mu^{*}\left(d_{\uparrow}\rightarrow\right)+\frac{1}{9}\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\cos^{2}\theta+\left(\frac{10}{9}\mu^{*}\left(d_{\uparrow}\rightarrow\right)+\frac{5}{9}\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\sin^{2}\theta$ (47)
$-\frac{8}{9}\left(\mu^{*}\left(d_{\uparrow}\rightarrow\right)-\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\sin\theta\cos\theta.$
To account for the contributions of orbital angular momentum polarizations ($\mu^{*L}_{B}$) towards the total baryonic magnetic moment, we have calculated the individual contributions from the valence quarks and quark sea as
$\mu_{B}^{*L}=\mu_{\mathrm{val}}^{*L}+\mu_{\mathrm{sea}}^{*L}.$ (48)
In this equation, $\mu_{\mathrm{val}}^{*L}$ and $\mu_{\mathrm{sea}}^{*L}$ represent the effective contributions of the valence and sea quarks towards the baryonic magnetic moment component, $\mu^{*L}_{B}$, due to the orbital angular momentum polarization. The details of the valence quark calculations has been discussed in ref. thirtyone . In the current section we will also describe the expressions used in the calculation of sea quark spin polarization in a baryon, which are obtained by substituting the third component of the orbital angular momentum of sea quarks in place of each valence quark as,
$q^{(\pm 1)}\rightarrow-T_{[q,GB]}q^{(\pm 1)}+\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}.\hskip 150.79968pt$ (49)
In eq. 49, $T_{[q,GB]}$ refers to the emission probability of a goldstone bosons (GB) from a quark $q^{(\pm 1)}$ and $\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}$ represents the transformation probability of a $q^{(\pm 1)}$ quark into two other quarks of $SU(3)$ model, given in thirtyone as
$\left|\psi\left(u^{(\pm 1)}\right)\right|^{2}=ad^{(\pm 1)}+a\alpha^{2}s^{(\pm 1)},$ (50)
$\left|\psi\left(d^{(\pm 1)}\right)\right|^{2}=au^{(\pm 1)}+a\alpha^{2}s^{(\pm 1)},$
$\left|\psi\left(s^{(\pm 1)}\right)\right|^{2}=a\alpha^{2}u^{(\pm 1)}+a\alpha^{2}d^{(\pm 1)}.$
The probability of transition a $q^{\pm 1}$ constituent quark into all the possible quarks ($u,d$ and $s$) along with GBs $\left(q\bar{q}^{\prime}\right)$ can be expressed as fiftyeight
$P\left(u^{\pm}\rightarrow u\bar{q}^{\prime}+q^{\prime\mp}\right)=\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)u^{\mp}+ad^{\mp}+a\alpha^{2}s^{\mp},$ (51)
$P\left(d^{\pm}\rightarrow d\bar{q}^{\prime}+q^{\prime\mp}\right)=au^{\mp}+\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)d^{\mp}+a\alpha^{2}s^{\mp},$
$P\left(s^{\pm}\rightarrow s\bar{q}^{\prime}+q^{\prime\mp}\right)=a\alpha^{2}u^{\mp}+a\alpha^{2}d^{\mp}+\frac{a}{3}\left(2\beta^{2}+\zeta^{\prime 2}\right)s^{\mp}.$
Using eq. 51, we can easily obtain the $T_{[q,GB]}$ for each constituent quark. The valence and sea quark contributions of the orbital angular momentum part of the magnetic moment of the baryons, can thus be calculated easily. As an example, we have presented the orbital angular momentum contribution to baryonic magnetic moment of $\frac{1}{2}^{-}$ $n^{*}$ resonances as,
$\mu^{*L}_{\mathrm{valence}}\left(n^{*}\right)=$ $\left(\frac{4}{9}\mu^{*}_{d}+\frac{2}{9}\mu^{*}_{u}\right)\cos^{2}\theta-\left(\frac{1}{9}\mu^{*}_{d}+\frac{2}{9}\mu^{*}_{u}\right)\sin^{2}\theta,$ (52)
$\mu^{*L}_{\mathrm{sea}}\left(n^{*}\right)=$ $\frac{2a}{9}\left[-\left(1+2\alpha^{2}\right)\mu^{*}_{d}+\left(1-\alpha^{2}\right)\mu^{*}_{u}\right.\left.+3\alpha^{2}\mu^{*}_{s}\right]\cos^{2}\theta$
$+\frac{a}{9}\left[-\left(1-\alpha^{2}\right)\mu^{*}_{d}\right.\left.+\left(1+2\alpha^{2}\right)\mu^{*}_{u}-3\alpha^{2}\mu^{*}_{s}\right]\sin^{2}\theta.$
Using similar methodology as demonstrated above, the magnetic moments of the $\frac{1}{2}^{-}$ octet baryon resonances $p^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$ and $\Xi^{*0}$ as well as the low-lying $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$, and $S_{11}^{0}(1650)$ have been calculated within the $\chi\mathrm{CQM}$ framework.
IV Results & Discussion
We now discuss numerical results on the in-medium masses (section IV.1) and magnetic moments (section IV.2) of spin $\frac{1}{2}^{-}$ octet baryon resonances. The effective baryon masses have been listed in tables 1 and 2, so as to facilitate a discussion on the mass modifications. As an outcome of the calculations discussed in the previous sections, the vacuum and in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances have been illustrated in tables 3, 4, 5 and 6. For the sake of comparison with the baryonic magnetic moments obtained from earlier studies, we have tabulated the values calculated through different models in table 7 for $\rho_{B}=0$.
IV.1 In-medium masses of $\frac{1}{2}^{-}$ baryon resonances
Figure 1: Effective masses of constituent quarks $u,d$ and $s$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
As the medium modifications of baryon masses is calculated using the medium modified masses of constituent quarks, we first discuss the impact of density and temperature on effective masses of quarks. In fig. 1, the in-medium masses, $m^{*}_{u},m^{*}_{d}$ and $m^{*}_{s}$, of the constituent quarks are shown as a function of density in units of nuclear saturation density, at $T=0$ MeV (on the left) and $T=100$ MeV (on the right)(where nuclear saturation density $\rho_{0}=0.16$ $\text{fm}^{-3}$). It also illustrates the in-medium modification in quark masses for different values of isospin asymmetry parameter ($i_{a}$) and strangeness fraction ($f_{s}$).
From fig. 1 it is observed that as we move from vacuum ($\rho_{B}=0$) to denser hadronic medium, the quark masses decrease monotonously. For symmetric nuclear matter at temperature $T=0$ MeV, the effective masses of light quarks decreases by 71.14% and that of $s$ quark decreases by 14.65%, at density value of $\rho_{B}=3\rho_{0}$. As observed, the effects of density on the $s$ quark is significantly lower than that on the light quarks, which is an implication to the negligible coupling of $s$ quark with $\sigma$ field fiftyfour ; sixtytwo . By introducing finite strangeness in the symmetric nuclear medium, at $f_{s}=0.3$ and $T=0$ MeV, with an increase in baryon density there is an appreciable drop in $m^{*}_{s}$, relative to that in symmetric nuclear matter, while in the light quarks only slight variations in the $m^{*}_{u}$ and $m^{*}_{d}$ values are observed. It is also observed that with the rise in temperature at $T=100$ MeV the effective masses of $u$ and $d$ quarks rises by 8.91 % at $\rho_{B}=\rho_{0}$ and the $s$ quark mass elevates by just 1.38 %. Thus an increase in the medium temperature causes a rise in the masses of constituent quarks, at the same density. The increase in isospin asymmetry of the medium leads to increase in-medium masses of constituent quarks. A study on the basis of an extended NJL model has previously demonstrated that while the in-medium masses of light quarks approach their current quark masses, the $s$ quark condensate remains large even at densities close to $\rho_{B}\approx 1$ $\text{fm}^{-3}$ xia .
In figs. 2, 3 and 4 and fig. 5 we have shown the variations in effective baryonic masses, as a function of medium density, for $\frac{1}{2}^{-}$ octet baryon resonances and $N^{*}$ resonance $S^{+(0)}_{11}(1650)$, respectively. Using eq. 25 the effective masses of $\frac{1}{2}^{-}$ baryon resonances are calculated in the isospin asymmetric strange matter, where the $E_{ispin}$ values are best fitted to procure the masses of $\frac{1}{2}^{-}$ baryon resonances in free space. A quick glance at the arguments and figures depicts, a monotonous decrease in the in-medium baryon masses as a function of density. The baryon masses in the medium fall rapidly in the lower density region up to $\rho_{B}=\rho_{0}$ and then saturate as the medium approaches large densities. As seen from fig. 2, when symmetric nuclear matter density rises upto $\rho_{B}=3\rho_{0}$, the mass of $p^{*}$ resonance drops appreciably by 33.2%. At a density of $3\rho_{0}$, with finite isospin asymmetry $i_{a}=0.3$, the mass of $p^{*}$ is 1010.03 MeV which is a rise from 1003.60 MeV at $i_{a}=0$. In fact, with higher valence $d$ quark content in a baryon, the density effects are found to be suppressed. This can be seen from the 25.27% decrease in the effective mass of $n^{*}$, as the medium density reaches $\rho_{B}=3\rho_{0}$ from vacuum. Keeping the density unchanged, say at $\rho_{B}=3\rho_{0}$, the $M^{*}_{n}$ value is found to incur a marginal increase, as a function of asymmetry of the medium, to attain 1149.18(1147.02) MeV for $i_{a}=0.3(0)$. In contrary to light quark content baryons, for symmetric nuclear matter the mass modifications incurred by $\Sigma^{*}$ resonances, as seen in fig. 3, is 18.5% and that of $\Xi^{*}$ resonances in fig. 4 is 14.5%. We can clearly state that, the density dependency on the baryons constituted by light quarks is much higher than that of hyperons, which is because $m^{*}_{u}$ and $m^{*}_{d}$ are more sensitive to density changes as compared to $m^{*}_{s}$. Additionally in hyperons, the in-medium effects are subject to the extent of $s$ quarks constituting a baryon. This is because as the number of constituent $s$ quarks increase, the in-medium effects on the constituent quark masses start having a dominating influence on the total mass of the baryon sixtytwo . In fig. 5, the masses of $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ varying as a function of medium density, have been shown. The extent of in-medium variations of the $N^{*}$ resonances is large compared to hyperons, like, in the case of $p^{*}$ and $n^{*}$ resonances. At a given value of $\rho_{B}$, with $i_{a}=0,T=0$ MeV, the impact of strangeness fraction, $f_{s}$, causes a sizable decrease in the masses of baryons. Looking up the mass of $\Xi^{*-}$ from table 1, at $\rho_{B}=3\rho_{0},f_{s}=0.3(0)$, we observe that the value decreases to 1399.14(1439.75) MeV. A cursory look at the plots of effective baryonic masses versus medium density suggests that, if the temperature of the medium increases to $T=100$ MeV then the mass modification curves will have relatively less steeper slope than that at $T=0$ MeV. For example, at $i_{a}=0,f_{s}=0(0.3)$ when the medium density is at $\rho_{B}=3\rho_{0}$, the effective mass of $\Sigma^{*+}$ being of 1300.88(1276.30) MeV, in table 1, at $T=0$ MeV increases to 1322.58(1287.69) MeV, in table 2, at $T=100$ MeV.
Owing to the strong coupling of $N^{*}$(1535) nucleon resonance with $\eta N$, there has been an earlier work that leverages $\eta$ mesic nuclei as a tool to map the reduction in-medium mass difference between nucleon $N$ and its chiral partner $N^{*}$, in chiral doublet model 2002_mass . In another study conducted using parity doublet model at finite baryonic densities of cold nuclear matter, the authors showcases the dependency of the vacuum mass of $N^{*}$ to determine the critical density and the order of the phase transition 2006_mass . For negative parity nucleons, the density effects on the in-medium baryonic mass turns out to be weaker than the impact on nucleon mass 2016_mass , which also gets reflected in fig. 2. In the context of hot hadronic matter, a study based on hadron resonance gas (HRG) sheds light on the temperature dependency of the effective masses of negative parity baryons 2018_mass . The observed temperature effects within $T=100$ MeV, have been in good agreement with the findings in current work. In the context of $\frac{1}{2}^{+}$ octet baryons, there has been studies on their in-medium mass modifications based on QMC hadron_mass , modified QMC (MQMC) mqmc , CQMF sixtytwo and also for decuplet, low-lying strange, charm, and bottom baryons in QMC framework tsushima1 ; tsushima2 .
Figure 2: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 3: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 4: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 5: Effective masses of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
IV.2 In-medium magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances
Figure 6: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 7: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 8: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 9: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 10: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 11: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 12: Effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
Figure 13: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV. Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
In this section we elaborately discuss the results of the vacuum and in-medium magnetic moments of ${\frac{1}{2}}^{-}$ octet baryon resonances $p^{*},n^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$, $\Xi^{*0}$ and low-lying $N^{*}$ resonances ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$, observed from figs. 6, 7, 8, 9, 10, 11, 12 and 13. For the sake of discussion, we have also included the individual contributions from valence and sea quarks to the total baryonic magnetic moment which allows us to better interpret the results.
In fig. 6, we have shown the magnetic moment of $p^{*}$ and $n^{*}$ resonances with density of the medium at temperatures $T=0$ MeV and $T=100$ MeV. The observed in-medium behavior of $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$, shows a steeper slope at lower densities of nuclear matter ($\rho_{B}\leq 2\rho_{0}$) and gradually attains saturation at much higher densities. As we trace the plot of these effective baryon magnetic moments while the medium density rises to $\rho_{B}=3\rho_{0}$, we observe that $\mu^{*}_{p^{*}}$ increases by $0.59\mu_{N}$ and $\mu^{*}_{n^{*}}$ initially approaches a minima at $-1.382\mu_{N}$ (at $\rho_{B}\approx\rho_{0}$) and then climbs up close to its vacuum value at $-1.375\mu_{N}$. A conclusive understanding of the observed effective baryonic magnetic moments of $p^{*}$ and $n^{*}$ can be provided by analyzing fig. 7, where the contributions from valence quarks, sea quarks and orbital angular momentum of the quark sea have been shown individually. For example, in case of $n^{*}$ in symmetric nuclear matter, the decrease in the valence quark contributions dominates over the increasing contributions of sea quarks and the orbital contributions of sea quarks, which leads to an overall decreasing behavior of $\mu^{*}_{n^{*}}$ below the saturation density of the medium ($\rho_{0}$). As the medium density moves past this point ($\rho_{B}\approx\rho_{0}$), owing to the dominance of $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$ contributions the effective in-medium behavior of $\mu^{*}_{n^{*}}$ reverses and it approaches higher values. On a similar note, the density dependency of $\mu^{*}_{p^{*}}$ can be comprehended by taking into account the increase in $\mu^{*}_{val}$ dominating over the decrease in $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$. In symmetric strange matter at $f_{s}=0.3$, the in-medium behavior of $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$ is enhanced. From fig. 6 we can infer that, the finite strangeness fraction in the medium leads to a splitting in baryonic magnetic moments of $p^{*}$ and $n^{*}$, which is visible beyond the nuclear saturation density, $\rho_{0}$. All other medium parameters affixed, non-zero isospin asymmetry ($i_{a}=0,0.3,0.5$) causes fine splitting in the $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$ curves as can be seen from fig. 6, which has also been identified in their respective $\mu^{*}_{val}$, $\mu^{*}_{sea}$, $\mu^{*}_{orbit}$ contributions by referring to fig. 7. As we have included “Goldstone Boson Exchange”(GBE) effect in this study, it is important to note that the sea quark effect has an opposite sign as compared to the valence quark effect, whereas the orbital angular momentum of sea quarks has the same contribution as that from the valence quark effect (see Table 3). The sign of these contributions can be justified, if we realize that the production of the quark sea is associated with a spin flip process (see eq. 30) and hence, the sea quarks has to have an opposite effect. An earlier study highlights the impact of Goldstone Boson (GB) degrees in chiral quark model ($\chi$QM), where the magnetic moments calculated for $p^{*}$ and $n^{*}$ in vacuum, by considering GBE interaction, are 1.4 $\mu_{N}$ and $-0.9\mu_{N}$ for $\chi$QM, which are slightly lower than the corresponding values, 1.7 $\mu_{N}$ and $-1.1\mu_{N}$ as obtained from CQM framework liu . In addition, the magnetic moments procured from $\chi$QM by considering Goldstone Boson Exchange (GBE) and One Gluon Exchange (OGE), separately, gives similar magnetic moment values and are in close agreement to our vacuum state results liu . Two other studies in the literature, with quasibound state approach quasibound and chiral unitary model chiral_unitary have previously calculated for magnetic moments of these two resonances and could be compared from table 7. Table 7 facilitates a comparitive study of the magnetic moments of $p^{*}$ and $n^{*}$ resonances obtained from various models.
Figure 8 depicts the in-medium modifications of $\mu^{*}_{\Sigma^{*+}}$, $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$, as a function of $\rho_{B}/\rho_{0}$ and the behavior of the individual contributions have been plotted in fig. 9. In nuclear medium, the magnetic moments of $\Sigma^{*+}$ initially increases up to $\rho_{B}=2\rho_{0}$ and then saturates at higher densities of the medium. By taking into account the individual contributions from fig. 9, we see that $\mu^{*}_{val}$ increases with density and it has a dominating effect over the cumulative depletion of $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$ contributions. In case of $\Sigma^{*-}$ (at $i_{a}=0,f_{s}=0$), the effective magnetic moment $\mu^{*}_{\Sigma^{*-}}$ decreases and gradually approaches a saturation at very high densities of $\rho_{B}\approx 5\rho_{0}$. Although the in-medium behavior of $\mu^{*}_{val}$ is accompanied by $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$, which imparts opposite contributions, but the valence quarks have a predominant role in determining the nature of $\mu^{*}_{\Sigma^{*-}}$. These opposite contributions due to sea quarks and the orbital motion of sea quarks at higher density regime, evidently explains the stagnant in-medium nature of $\mu^{*}_{\Sigma^{*-}}$ at large medium densities. Further, the density dependent medium modification of $\mu^{*}_{\Sigma^{*0}}$, in symmetric nuclear matter, is much alike the in-medium nature of $\mu^{*}_{\Sigma^{*+}}$ except for a slightly decreasing slope at densities above $1.5\rho_{0}$. The dominance of $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions over the $\mu^{*}_{sea}$ contribution, justifies the in-medium curve of $\mu^{*}_{\Sigma^{*0}}$ in a nuclear medium. From a point of comparison if you see Table 7, the magnetic moments of $\mu^{*}_{\Sigma^{*+}}$ (1.647 $\mu_{N}$), $\mu^{*}_{\Sigma^{*-}}$ $(-0.754\mu_{N})$ and $\mu^{*}_{\Sigma^{*0}}$ (0.446 $\mu_{N}$) in free space are found to be in agreement with the NCQM results 1.814 $\mu_{N}$, $-0.689\mu_{N}$ and 0.820 $\mu_{N}$, respectively. As seen from Tables 3 and 5, for $f_{s}=0(0.3)$ at $\rho_{B}=3\rho_{0}$, $\mu^{*}_{\Sigma^{*+}}$ and $\mu^{*}_{\Sigma^{*-}}$ values attain 1.697(1.723) $\mu_{N}$ and $-0.770(-0.779)\mu_{N}$. This suggests that as we introduce finite strangeness fraction to the non-strange medium, there will be an enhancement in the nature of in-medium variations of baryonic magnetic moments of $\Sigma^{*+}$, $\Sigma^{*0}$ and $\Sigma^{*-}$. For finite $f_{s}$, the magnetic moments of $\Sigma^{*+}$ and $\Sigma^{*0}$ surges while that of $\Sigma^{*-}$ approaches lower values, instead of attaining a saturation at large densities. For an instance, the in-medium values of $\mu^{*}_{\Sigma^{*+}}$, $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$ (in units of $\mu_{N}$) at $i_{a}=0(0.5)$, when referred to Tables 3 and 4, are $1.697(1.698),-0.770(-0.763)and0.463(0.472)$ in a medium density of $3\rho_{0}$. This and a glance into fig. 8 makes it apparent that the impact of non-zero $i_{a}$, a fine splitting in effective baryonic magnetic moments, is more visible for $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$ as compared to $\mu^{*}_{\Sigma^{*+}}$. Such $i_{a}$ dependent fine shifts in the baryonic magnetic moment have been previously found in works, in the present formalism, dedicated to octet baryons fiftythree and decuplet baryons fiftyfour .
We find it interesting to note that for $\Xi^{*}$ baryons the density effects on $\mu^{*}_{\Xi^{*-}}$ and $\mu^{*}_{\Xi^{*0}}$ values show impressive similarity, owing to the valence quark spin polarizations and corresponding orbital moments of two $s$ quarks that suppresses the impact of light $u$ and $d$ quarks in $\Xi^{*-}$ and $\Xi^{*0}$, respectively sixtytwo . The baryonic magnetic moment of $\Xi^{*-}$, as seen from fig. 10, rises steeply with increase in the relative density ($\rho_{B}/\rho_{0}$) of the medium. Figure 11 clearly illustrates that for $\Xi^{*-}$, the cumulative increasing behavior of $\mu^{*}_{val}$ and $\mu^{*}_{sea}$ contributions overshadows the decreasing orbital contributions from sea quarks ($\mu^{*}_{orbit}$). In a similar manner, a quantitative analysis of fig. 10 and fig. 11 suggests that the density dependent increase in $\mu^{*}_{\Xi^{*0}}$ is well justified by the contributions coming from $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$, which suppresses the contributions coming from $\mu^{*}_{sea}$. Much alike other baryons, we have observed a correlation between finite isospin asymmetry of the medium and a distinguishable splitting in the baryonic magnetic moments in case of $\Xi^{*}$ resonances. Now in order to understand the temperature effects on the in-medium baryonic magnetic moments, we shall lay emphasis on the case of $\mu^{*}_{\Xi^{*-}}$. A close look at fig. 10, shows that for symmetric nuclear matter at medium densities of $\rho_{0}$ and $3\rho_{0}$, at $T=0(100)$ MeV the $\mu^{*}_{\Xi^{*-}}$ values obtained, in $\mu_{N}$ units, are 0.313(0.286) and 0.179(0.163), respectively. This reveals the in-medium nature of $\mu^{*}_{\Xi^{*-}}$, an increasing curve as a function of medium density, is subject to slight reduction in its value as the temperature rises to $T=100$ MeV. In addition, the vacuum state magnetic moments of $\Xi^{*-}$ resonances are found to be much lower than those procured from NCQM (see Table 7).
From fig. 12, we observe that the magnetic moments of ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$ increases, overall as a function of density of the hadronic matter $\rho_{B}$ (in $\rho_{0}$ units). As from fig. 13 it is seen that, the $\mu^{*}_{sea}$ exercises control over the $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions, which explains the medium modifications of the magnetic moment of ${S^{+}_{11}}(1650)$ in lower density regime (i.e. up till $\rho_{B}=2\rho_{0}$). However, as the nuclear medium grows denser the $\mu^{*}_{val}$ and $\mu^{*}_{orbit}$ contributions demonstrates a dominating influence on the total effective magnetic moment of ${S^{+}_{11}}(1650)$. From fig. 13 it is also observed that for the $N^{*}$ resonance ${S^{0}_{11}}(1650)$, the value of $\mu_{{S^{0}_{11}}(1650)}$ receives major contributions, monotonously, due to the dominant valence quark contribution ($\mu^{*}_{val}$) decreasing with rise in density. For a comparative study based on different models, similar justifications as discussed in the case of $p^{*}$ and $n^{*}$ resonances suffices and the results have been tabulated in Table 7.
A study based on QMC as well as MQMC, for octet baryons, shows that we can expect sizable influence on effective baryonic moments due to temperature above $T=150$ MeV at lower medium densities mqmc . Speaking of correlations, the authors in ref. seventyfour have discussed a strong correlation of the baryonic magnetic moment with that of the bag radius of a baryon, as the two quantities have a similar behavior as density dependent functions. Since this is the first work on medium modifications of $\frac{1}{2}^{-}$ baryon resonances, so referring to previous works on in-medium density effects on magnetic moments of baryons using the current formalism sixtytwo ; fiftyfour and QMC tsushima2 , might facilitate the purpose of a comparative study for future work on the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances.
V Summary
To summarize, in the present work we have investigated the modifications induced on the masses as well as the magnetic moments of $J_{p}={\frac{1}{2}}^{-}$ baryon resonances in isospin asymmetric strange hadronic matter. In light of chiral $SU(3)$ quark mean field model the in-medium effects are introduced and we computed the effective baryon masses by using the masses of constituent quarks. The mass adjusted quark magnetic moments are then calculated, to include relativistic quark confinement, by using the baryon masses as inputs. In hadronic medium the in-medium masses of lights quarks, $u$ $d$, vary significantly in comparison to the observed changes in medium modified masses of $s$ quark. In the $\chi$CQM formalism, we then computed the baryonic magnetic moments from the contributions due to valence quarks, sea quarks and orbital angular momentum due of sea quarks. As derived from the results is that, the medium modifications are more prominent for the baryons comprised of light $u$ or $d$ quarks, than for hyperons. For strange matter, the in-medium constituent quark masses, effective baryonic masses as well as the effective baryonic magnetic moments suffer appreciable enhancement in their in-medium behaviour. By carrying out this study at finite temperatures and finite $i_{a}$ values of the hadronic medium, we have noted a fine shift in the in-medium baryonic magnetic moments, although the temperature effects have been negligible below $T=150$ MeV.
Baryon Vacuum Mass $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$f_{s}=0$ $f_{s}=0.3$ $f_{s}=0$ $f_{s}=0.3$
(MeV) $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$
$M_{S^{+}_{11}(1650)}^{*}$ 1650.0 $1439.01$ $1442.02$ $1435.09$ $1440.78$ $1263.40$ $1268.27$ $1254.90$ $1260.31$
$M_{S^{0}_{11}(1650)}^{*}$ 1650.0 $1439.01$ $1439.98$ $1435.09$ $1438.69$ $1263.40$ $1265.55$ $1254.90$ $1257.48$
$M_{p^{*}}^{*}$ 1535.0 $1241.14$ $1245.05$ $1225.47$ $1233.20$ $1003.60$ $1010.03$ $953.84$ $962.62$
$M_{n^{*}}^{*}$ 1535.0 $1323.31$ $1324.28$ $1319.37$ $1322.98$ $1147.02$ $1149.18$ $1138.50$ $1141.08$
$M_{\Sigma^{*+}}^{*}$ 1620.0 $1442.84$ $1446.59$ $1434.69$ $1440.74$ $1300.88$ $1306.54$ $1276.30$ $1283.18$
$M_{\Sigma^{*-}}^{*}$ 1620.0 $1442.84$ $1442.48$ $1434.69$ $1436.54$ $1300.88$ $1301.06$ $1276.30$ $1277.49$
$M_{\Sigma^{*0}}^{*}$ 1620.0 $1442.84$ $1444.53$ $1434.69$ $1438.64$ $1300.88$ $1303.80$ $1276.30$ $1280.34$
$M_{\Xi^{*-}}^{*}$ 1690.0 $1547.34$ $1547.70$ $1534.99$ $1537.17$ $1439.75$ $1440.69$ $1399.14$ $1401.79$
$M_{\Xi^{*0}}^{*}$ 1690.0 $1547.34$ $1549.75$ $1534.99$ $1539.27$ $1439.75$ $1443.43$ $1399.14$ $1404.64$
[…]
Baryon Vacuum Mass $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$f_{s}=0$ $f_{s}=0.3$ $f_{s}=0$ $f_{s}=0.3$
(MeV) $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$ $i_{a}=0$ $i_{a}=0.3$
$M_{S^{+}_{11}(1650)}^{*}$ 1650.0 $1469.66$ $1471.49$ $1454.68$ $1455.95$ $1289.10$ $1292.78$ $1270.06$ $1272.66$
$M_{S^{0}_{11}(1650)}^{*}$ 1650.0 $1469.66$ $1469.70$ $1454.68$ $1453.87$ $1289.10$ $1289.73$ $1270.06$ $1269.38$
$M_{p^{*}}^{*}$ 1535.0 $1283.87$ $1286.17$ $1251.81$ $1253.30$ $1040.65$ $1045.23$ $972.85$ $976.01$
$M_{n^{*}}^{*}$ 1535.0 $1354.07$ $1354.10$ $1339.04$ $1338.22$ $1172.82$ $1173.45$ $1153.71$ $1153.02$
$M_{\Sigma^{*+}}^{*}$ 1620.0 $1468.45$ $1471.06$ $1450.57$ $1452.89$ $1322.58$ $1327.39$ $1287.69$ $1291.80$
$M_{\Sigma^{*-}}^{*}$ 1620.0 $1468.45$ $1467.46$ $1450.57$ $1448.72$ $1322.58$ $1321.26$ $1287.69$ $1285.20$
$M_{\Sigma^{*0}}^{*}$ 1620.0 $1468.45$ $1469.26$ $1450.57$ $1450.81$ $1322.58$ $1324.33$ $1287.69$ $1288.50$
$M_{\Xi^{*-}}^{*}$ 1690.0 $1567.79$ $1567.58$ $1547.08$ $1546.28$ $1457.33$ $1457.12$ $1406.70$ $1405.71$
$M_{\Xi^{*0}}^{*}$ 1690.0 $1567.79$ $1569.38$ $1547.08$ $1548.36$ $1457.33$ $1460.19$ $1406.70$ $1409.02$
[…]
Baryons $\rho_{B}=0$ $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$
${S^{+}_{11}(1650)}\left(\mu_{N}\right)$ 0.140 -0.169 -0.136 -0.165 0.161 -0.200 -0.051 -0.090 0.184 -0.236 0.061 0.008
${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.065 -0.286 0.125 0.904 1.213 -0.334 0.064 0.943
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.299 -0.392 0.207 2.114 2.843 -0.483 0.016 2.376
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.464 0.263 -0.181 -1.382 -1.689 0.305 0.007 -1.376
${\Sigma}^{*+}\left(\mu_{N}\right)$ 1.761 -0.343 0.229 1.647 1.935 -0.382 0.132 1.684 2.086 -0.419 0.030 1.697
${\Sigma}^{*-}\left(\mu_{N}\right)$0.694 0.133 -0.193 -0.754 -0.821 0.153 -0.091 -0.760 -0.971 0.174 0.026 -0.770
${\Sigma}^{*0}\left(\mu_{N}\right)$ 0.533 -0.105 0.017 0.446 0.556 -0.114 0.020 0.461 0.557 -0.122 0.028 0.463
${\Xi}^{*-}\left(\mu_{N}\right)$0.237 0.221 0.116 0.100 -0.106 0.231 0.052 0.177 0.093 0.233 -0.017 0.309
${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.833 -0.282 -0.084 -1.200 -0.687 -0.318 -0.019 -1.026
Table 3: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric nuclear matter ($i_{a}=0$ and $f_{s}=0$) at a medium temperature $T=0$ MeV.
Baryons $\rho_{B}=0$ $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$
${S^{+}_{11}(1650)}\left(\mu_{N}\right)$ 0.140 -0.169 -0.136 -0.165 0.160 -0.199 -0.047 -0.085 0.182 -0.233 0.064 0.013
${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.062 -0.285 0.134 0.911 1.205 -0.331 0.079 0.953
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.280 -0.389 0.213 2.104 2.792 -0.475 0.032 2.349
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.459 0.263 -0.185 -1.382 -1.676 0.303 -0.003 -1.376
${\Sigma}^{*+}\left(\mu_{N}\right)$ 1.761 -0.343 0.229 1.647 1.927 -0.380 0.137 1.684 2.072 -0.416 0.042 1.698
${\Sigma}^{*-}\left(\mu_{N}\right)$0.694 0.133 -0.193 -0.754 -0.818 0.152 -0.092 -0.757 -0.960 0.173 0.024 -0.763
${\Sigma}^{*0}\left(\mu_{N}\right)$ 0.533 -0.105 0.017 0.446 0.557 -0.114 0.023 0.466 0.561 -0.122 0.033 0.472
${\Xi}^{*-}\left(\mu_{N}\right)$0.237 0.221 0.116 0.100 -0.116 0.231 0.052 0.168 0.065 0.234 -0.015 0.284
${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.840 -0.281 -0.087 -1.209 -0.710 -0.314 -0.026 -1.051
Table 4: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric nuclear matter ($i_{a}=0.5$ and $f_{s}=0$) at a medium temperature $T=0$ MeV.
Baryons $\rho_{B}=0$ $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$
${S^{+}_{11}(1650)}\left(\mu_{N}\right)$ 0.140 -0.169 -0.136 -0.165 0.162 -0.200 -0.049 -0.088 0.185 -0.238 0.068 0.015
${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.068 -0.287 0.124 0.904 1.221 -0.337 0.062 0.946
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.328 -0.397 0.205 2.136 2.991 -0.509 0.005 2.487
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.468 0.264 -0.178 -1.382 -1.701 0.308 0.018 -1.375
${\Sigma}^{*+}\left(\mu_{N}\right)$ 1.761 -0.343 0.229 1.647 1.946 -0.384 0.130 1.692 2.125 -0.427 0.026 1.723
${\Sigma}^{*-}\left(\mu_{N}\right)$0.694 0.133 -0.193 -0.754 -0.826 0.153 -0.089 -0.762 -0.991 0.177 0.034 -0.779
${\Sigma}^{*0}\left(\mu_{N}\right)$ 0.533 -0.105 0.017 0.446 0.560 -0.115 0.020 0.464 0.567 -0.125 0.030 0.472
${\Xi}^{*-}\left(\mu_{N}\right)$0.237 0.221 0.116 0.100 -0.108 0.233 0.051 0.176 0.100 0.239 -0.021 0.319
${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.841 -0.285 -0.083 -1.209 -0.703 -0.328 -0.016 -1.048
Table 5: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric hyperonic matter ($i_{a}=0$ and $f_{s}=0.3$) at a medium temperature $T=0$ MeV.
Baryons $\rho_{B}=0$ $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
$\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$ $\mu^{*}_{\text{val}}$ $\mu^{*}_{\text{sea}}$ $\mu^{*}_{\text{orbit}}$ $\mu_{B}^{*}$
${S^{+}_{11}(1650)}\left(\mu_{N}\right)$ 0.140 -0.169 -0.136 -0.165 0.160 -0.198 -0.047 -0.085 0.183 -0.235 0.074 0.022
${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.061 -0.284 0.135 0.911 1.216 -0.334 0.075 0.957
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.295 -0.392 0.216 2.119 2.944 -0.501 0.016 2.459
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.458 0.262 -0.187 -1.382 -1.694 0.306 0.012 -1.375
${\Sigma}^{*+}\left(\mu_{N}\right)$ 1.761 -0.343 0.229 1.647 1.933 -0.381 0.138 1.690 2.112 -0.424 0.035 1.723
${\Sigma}^{*-}\left(\mu_{N}\right)$0.694 0.133 -0.193 -0.754 -0.818 0.152 -0.093 -0.758 -0.982 0.176 0.034 -0.771
${\Sigma}^{*0}\left(\mu_{N}\right)$ 0.533 -0.105 0.017 0.446 0.560 -0.115 0.023 0.468 0.570 -0.124 0.034 0.480
${\Xi}^{*-}\left(\mu_{N}\right)$0.237 0.221 0.116 0.100 -0.122 0.233 0.053 0.163 0.076 0.240 -0.021 0.295
${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.851 -0.282 -0.088 -1.222 -0.722 -0.324 -0.022 -1.069
Table 6: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric hyperonic matter ($i_{a}=0.5$ and $f_{s}=0.3$) at a medium temperature $T=0$ MeV.
Baryons $\chi$CQM $\chi$QMliu CQMliu NCQMthirtyone Quasiboundquasibound CUchiral_unitary
$\mu_{\text{val}}$ $\mu_{\text{sea}}$ $\mu_{\text{orbit}}$ $\mu_{B}$ $\mu_{B}$ (GBE) $\mu_{B}$ (OGE) $\mu_{B}$ (GBE) $\mu_{B}$ $\mu_{B}$ $\mu_{B}$
${S^{+}_{11}(1650)}\left(\mu_{N}\right)$ 0.140 -0.169 -0.136 -0.165 0.1 0.0 0.3 0.106 - -
${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 0.6 0.7 0.8 0.951 - -
${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 1.4 1.6 1.7 1.894 1.86 1.13
${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -0.9 -1.0 -1.1 -1.284 -0.56 -0.248
${\Sigma}^{*+}\left(\mu_{N}\right)$ 1.761 -0.343 0.229 1.647 - - - 1.814 - -
${\Sigma}^{*-}\left(\mu_{N}\right)$0.694 0.133 -0.193 -0.754 - - - -0.689 - -
${\Sigma}^{*0}\left(\mu_{N}\right)$ 0.533 -0.105 0.017 0.446 - - - 0.820 - -
${\Xi}^{*-}\left(\mu_{N}\right)$0.237 0.221 0.116 0.100 - - - -0.315 - -
${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 - - - -0.990 - -
Table 7: The above table allows a comparative study of magnetic moments of $\frac{1}{2}^{-}$ baryon resonances, computed using different models in the free space ($\rho_{B}=0$).
VI APPENDIX
By minimizing the thermopotential in eq. 27, we derive the equations of motion for scalar fields $\sigma$, $\zeta$, the dilaton field $\chi$, scalar iso-vector field $\delta$ and the vector fields $\omega,\rho,\phi$ in view of eq. 29 as in fiftyfour
[…]
$\,+\,\frac{\chi^{2}}{\chi_{0}^{2}}m_{\pi}^{2}f_{\pi}-\left(\frac{\chi}{\chi_{0}}\right)^{2}m_{\omega}\omega^{2}\frac{\partial m_{\omega}}{\partial\sigma}\,-\left(\frac{\chi}{\chi_{0}}\right)^{2}m_{\rho}\rho^{2}\frac{\partial m_{\rho}}{\partial\sigma}\,+\,\sum_{i}\frac{\partial M_{i}^{\ast}}{\partial\sigma}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}$ (53)
[…]
$+\,\frac{\chi^{2}}{\chi_{0}^{2}}\left(\sqrt{2}m_{K}^{2}f_{K}\,-\,\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\right)+\,\sum_{i=\Lambda,\Sigma^{\pm,0},\Xi^{-,0}}\frac{\partial M_{i}^{\ast}}{\partial\zeta}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}=0,$ (54)
$\frac{\partial\Omega}{\partial\chi}=k_{0}\chi\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)-k_{3}\left(\sigma^{2}-\delta^{2}\right)\zeta+\frac{2\chi}{\chi_{0}^{2}}\left[m_{\pi}^{2}f_{\pi}\sigma+\left(\sqrt{2}m_{K}^{2}f_{K}\,-\,\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\right)\zeta\right]$
$-\frac{\chi}{\chi_{0}^{2}}\left(m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2}\right)+\left(4k_{4}+1+{\rm ln}\frac{\chi^{4}}{\chi_{0}^{4}}-\frac{4\xi}{3}{\rm ln}\left(\left(\frac{\left(\sigma^{2}-\delta^{2}\right)\zeta}{\sigma_{0}^{2}\zeta_{0}}\right)\right)\right)\chi^{3}=0,$ (55)
[…]
$-\,\sum_{i}g_{\delta i}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}$ (56)
$\frac{\partial\Omega}{\partial\omega}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\omega}^{2}\omega+4g_{4}\omega^{3}+12g_{4}\omega\rho^{2}-\sum_{i}g_{\omega i}\rho_{i}=0,$ (57)
$\frac{\partial\Omega}{\partial\rho}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\rho}^{2}\rho+4g_{4}\rho^{3}+12g_{4}\omega^{2}\rho-\sum_{i}{g_{\rho i}}\rho_{i}=0,$ (58)
and
$\frac{\partial\Omega}{\partial\phi}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\phi}^{2}\phi+8g_{4}\phi^{3}-\sum_{i}g_{\phi\,i}\rho_{i}=0.$ (59)
In eqs. 57, 58 and 59 $\rho_{i}$ represents the vector density of baryons and is given by,
$\rho_{i}=\gamma_{i}\int\frac{d^{3}k}{(2\pi)^{3}}\Big{(}f_{i}(k)-\bar{f}_{i}(k)\Big{)},$ (60)
and in sections VI, VI and VI $\rho_{i}^{s}$ stands for the scalar density of baryons, given as
$\rho_{i}^{s}=\gamma_{i}\int\frac{d^{3}k}{(2\pi)^{3}}\frac{M_{i}^{*}}{E^{\ast}_{i}(k)}\Big{(}f_{i}(k)+\bar{f}_{i}(k)\Big{)}.$ (61)
Here, $f_{i}(k)$ and $\bar{f}_{i}(k)$ appearing in eqs. 60 and 61 represent the Fermi distribution functions for fermions and anti-fermions, at finite temperature and are expressed as
$f_{i}(k)$ $=$ $\frac{1}{1+\exp\left[(E^{*}_{i}(k)-\nu^{*}_{i})/k_{B}T\right]}~{}~{}~{}\text{and}~{}~{}~{}\bar{f}_{i}(k)=\frac{1}{1+\exp\left[(E^{*}_{i}(k)+\nu^{*}_{i})/k_{B}T\right]}~{}.$ (62)

### v2

@@
-Magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances in hot and dense strange hadronic matter
+Magnetic moments of $\frac{1}{2}^{-}$ baryon resonances in strange hadronic matter at high densities
 Abstract
-This work comprises the calculations of magnetic moments of $\frac{1}{2}^{-}$ baryon resonances in the presence of hot and dense hadronic matter.
-Using the chiral $SU(3)$ quark mean field model, we have taken into account the impact on in-medium scalar meson fields to investigate the density effects on the in-medium baryon masses and their constituent quarks.
-In light of chiral constituent quark model $\chi$CQM, we have calculated the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances and scrutinized the explicit contributions coming from the valence quarks, sea quarks and the orbital angular momentum of sea quarks.
-The effective baryonic magnetic moments have been investigated for the varying values of isospin asymmetry and strangeness fraction.
-I Introduction
-In quantum chromodynamics (QCD), quarks and gluons are treated as fundamental degrees of freedom which form bound states known as hadrons (baryons($qqq$) and mesons($q\bar{q}$)).
+This work primarily focusses on determining the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances in the presence of hot and dense hadronic matter.
+In the chiral $SU(3)$ quark mean field model approach, we have essentially accounted for the effects on in-medium scalar meson fields to investigate the impact of high densities on the in-medium baryon masses and their constituent quarks.
+In light of chiral constituent quark model $\chi$CQM, we have calculated the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances and scrutinized the effects due to its internal constituents: the valence quarks, sea quarks and the orbital moment of sea quarks.
+Furthermore, we have investigated the effective baryonic magnetic moments for the finite magnitudes of isospin asymmetry and strangeness fraction.
+[…]
+xxxx, xxx
+1 Introduction
+The theory of quantum chromodynamics (QCD), is based on the consideration of quarks and gluons as the fundamental degrees of freedom which form bound states known as hadrons (baryons($qqq$) and mesons($q\bar{q}$)).
 These fundamental degrees of freedom undergo dynamical changes due to strong interactions in dense nuclear medium, which in turn cause in-medium modifications in the properties of hadrons.
 Confinement makes it challenging to understand QCD at lower energy scales.
-Scrutinizing the internal structure and the properties of light hadrons allow us to unravel the physics of strong interactions at low energies.
+Studying the internal structure and the properties of light hadrons allow us to unravel the physics of strong interactions at lower energy regimes.
 Apart from heavy-ion collision experimental studies on chiral symmetry restoration at high temperature thirtyseven , and the de-confinement of hadrons to quark-gluon plasma (QGP) thirtyeight , it has become a primary interest for experiments such as PANDA at FAIR thirtythree ; thirtyfour , BESIII at BEPCII thirtysix ; thirtytwo , HADES at GSI thirtyfive etc., to explore the electromagnetic properties of hadrons in dense matter.
@@
 Magnetic moment of hadrons, in free space, play a very important role in investigating their internal structure at a sub-nuclear level sixtyfour .
-Theoretically, the magnetic moments of octet as well as decuplet baryons have been extensively studied in free space one ; two ; three ; four ; five ; six ; seven ; eight ; nine .
+In theory, there has been extensive studies on the magnetic moments of octet and decuplet baryons in free space across a range of different models one ; two ; three ; four ; five ; six ; seven ; eight ; nine .
 Isgur framed his model with one-gluon exchange, in order to account for internal structure of hadrons isgur1 ; isgur2 ; isgur3 ; isgur4 .
-The constituent quark model (CQM) propounded that the baryonic magnetic moments can be calculated as the summation of the individual magnetic moments of its constituent quarks ten ; eleven .
-It was Glozman who incorporated the Goldstone boson exchange (GBE) interaction in the low energy regime and the concept of one-gluon exchange was substituted by vector meson exchange interactions, which in fact led to better fits for baryon spectra glozman .
+The constituent quark model (CQM) propounded that the baryonic magnetic moments are given as the summation of the individual magnetic moments of its constituent quarks ten ; eleven .
+It was Glozman who incorporated the Nambu Goldstone boson exchange (NGBE) interaction in the low energy regime and the concept of one-gluon exchange was substituted by vector meson exchange interactions, which in fact led to better fits for baryon spectra glozman .
 It was not until the development of chiral constituent quark model ($\chi\mathrm{CQM}$) that the pseudo-scalar and scalar meson exchanges were considered which was later extended to the inclusion of vector meson exchange dai .
@@
 In previous studies, this model has been used for calculating the baryonic magnetic moments, by taking into account the sea quark polarizations and their orbital angular momentum contribution through the Cheng-Li mechanism twentyeight ; twentynine ; thirty .
-In the $\chi$CQM formalism, the magnetic moments of low-lying ${\frac{1}{2}}^{-}$ $N^{*}$ resonances thirtyone , ${\frac{1}{2}}^{-}$ octet baryon resonances thirtyone and low-lying ${\frac{1}{2}}^{-}$ $\Lambda$ resonances A_Torres have been calculated in earlier studies.
+In the $\chi$CQM formalism, the magnetic moments of low-lying ${\frac{1}{2}}^{-}$ $N^{*}$ resonances thirtyone , ${\frac{1}{2}}^{-}$ octet baryon resonances thirtyone and low-lying ${\frac{1}{2}}^{-}$ $\Lambda$ resonances A_Torres have been calculated in free space, in earlier studies.
 Apart from $\chi$CQM, the magnetic moments of $N^{*}$ resonances have been computed by three quark Hamiltonian approach and the results were not consistent with lattice QCD values with very large variations hamiltonian_approach .
@@
 As mentioned earlier, the study of magnetic moments of baryons and the medium dependent effects on the electromagnetic properties is an area of active research and stands relevant to some of the experiments.
-In order to study the properties of nuclear matter, various theoretical models like the MIT bag model thirtynine , linear sigma model Petro , non-linear sigma model fortyfive , Zimanyi and Moszkowski model seventyone , cloudy bag model fortythree , Nambu-Jona-Lasinio (NJL) model fortytwo , light-front quark model Schlumpf1993 , parity doublet model paritydoublet1 ; paritydoublet2 etc., have been developed.
+For studying the properties of nuclear matter, various theoretical models like the MIT bag model thirtynine , linear sigma model Petro , non-linear sigma model fortyfive , Zimanyi and Moszkowski model seventyone , cloudy bag model fortythree , Nambu-Jona-Lasinio (NJL) model fortytwo , light-front quark model Schlumpf1993 , parity doublet model paritydoublet1 ; paritydoublet2 etc., have been developed.
 The cloudy bag model attained increased accuracy over MIT bag model in explaining QCD effects by incorporating the presence of pion clouds at the core of a baryon.
@@
 For octet baryons in light of the QMC model effoctet , it has been shown that the in-medium electric form factors are dampened and the magnetic form factors are enhanced, and this finding is consistent with the in-medium variations of the bound proton electromagnetic form factors observed at Jefferson Lab jlab .
-The study of effective magnetic moments of baryons was also extended to calculate the low-lying charm, and the low-lying bottom baryons with the aim of understanding the roles of the heavy and light quarks in symmetric nuclear matter tsushima1 ; tsushima2 .
+There has been studies on the effective baryonic magnetic moments aimed to carry out calculation for the low-lying charm and the low-lying bottom baryons, to examine the roles of the heavy and light quarks in symmetric nuclear matter tsushima1 ; tsushima2 .
 While a comparative study on octet baryons with modified quark meson coupling model (MQMC) mqmc , suggests that the observed in-medium baryonic magnetic moments incur drastic variations compared to that in QMC framework.
@@
 The CQMF model has recently been used to study valence quark distribution and transverse momentum dependent distributions of valence quarks of pions Puhan:2024xdq ; Kaur:2024wze .
-Presence of strangeness fraction ($f_{s}$) is a principal aspect to understand the departure of hadronic properties of baryons in strange hadronic matter comprising of hyperons from that in a non-strange medium.
-The in-medium variations of magnetic moments of octet and decuplet baryons in strange medium has shown stark deviation, as studied in fiftythree ; fiftyfour ; sixtytwo , for baryons of higher strangeness quantum number.
-In the present work, we have focused our study on the in-medium effects on the magnetic moments of the low-lying $N^{*}$ resonances and $\frac{1}{2}^{-}$ octet baryon resonances in CQMF model.
+In the current work we calculate the magnetic moments of the low-lying $N^{*}$ resonances and $\frac{1}{2}^{-}$ octet baryon resonances in hot and dense hadronic matter.
 It is exciting to study these resonances, as their detection connects hadron models to QCD properties in low and intermediate energy regimes.
@@
 The large decay widths of these excited baryon resonances, makes it challenging to detect them in experiments.
-In our current work, we have incorporated $\chi\mathrm{CQM}$ to calculate the properties of the baryon resonances as in thirtyone and have visualized them as a system of three valence quarks surrounded by the meson cloud.
+While Ref.
+thirtyone emphasizes on magnetic moment calculations of negative parity octet baryons in free space, our study extends beyond free space in hot and dense hadronic medium.
+As we shall see later, to calculate the magnetic moments of constituent quarks the impact of relativistic quark confinement is taken into account in the calculations of present work fiftyfive .
+However, in Ref.
+thirtyone such confinement effects have not been considered and this causes a difference in the values of vacuum baryonic magnetic moments of this study as compared to our current calculations.
+Presence of strangeness fraction ($f_{s}$) is a principal aspect to understand the departure of hadronic properties of baryons in strange hadronic matter comprising of hyperons, from that in a non-strange medium.
+The in-medium variations of magnetic moments of octet and decuplet baryons in strange medium has shown stark deviation, as studied in refs.
+fiftythree ; fiftyfour ; sixtytwo , for baryons of higher strangeness quantum number.
+The effects of increasing isospin asymmetry has also been considered for the analysis of density dependent in-medium modifications of baryons.
+This manuscript has been organized as follows: Section 2 comprises of two subsections describing the theoretical models used in the current work.
+In section 2.1 the CQMF model used to compute the in-medium quark masses and effective baryon masses, at finite temperature and density of the medium is presented.
+In section 2.2 we have calculated the contributions of valence quarks, sea quarks and orbital angular momentum of sea quarks towards the total magnetic moments of baryons, in light of $\chi\mathrm{CQM}$ model.
+Section 3 is dedicated to a detailed discussion on the results obtained from numerical calculations.
+We devote section 4 to present the summary of the current work.
+2 Theoretical Framework
+Central to this work, we have used $\chi\mathrm{CQM}$ to calculate the properties of the baryon resonances as in Ref.
+thirtyone and have visualized them as a system of three valence quarks.
 We have calculated the in-medium magnetic moments of the ${\frac{1}{2}}^{-}$ baryon resonances at finite temperature and density of hadronic medium through the medium modification of the corresponding constituent quark masses and the baryon masses.
@@
 The confined quarks inside the baryons, participate in the exchange interaction with these scalar meson fields as well as vector meson fields $\omega$, $\rho$ and $\phi$.
-This manuscript has been organized as follows: In section II the CQMF model used to compute the in-medium quark masses and effective baryon masses, at finite temperature and density of the medium is presented.
-In section III we have calculated the contributions of valence quarks, sea quarks and orbital angular momentum of sea quarks towards the total magnetic moments of baryons, in light of $\chi\mathrm{CQM}$ model.
-Section IV is dedicated to a detailed discussion on the results obtained from numerical calculations.
-We devote section V to present the summary of the current work.
-II Chiral Quark Mean Field Model
-In this section, we discuss briefly the framework, CQMF model, to calculate the in-medium constituent quark and baryon masses.
-As discussed earlier, in the CQMF model the quarks and mesons are treated as fundamental degrees of freedom, and the quarks are assumed to be confined inside baryons by an effective confining potential.
-The constituent quarks and mesons obtain their masses through the inclusion of properties such as spontaneous and explicit (for pseudoscalar mesons) symmetry breaking fortyfour .
+2.1 Chiral Quark Mean Field Model
+In this section, the CQMF model is discussed briefly which serves as the basic framework to calculate the in-medium masses of constituent quark and baryon.
+As discussed earlier, in CQMF model the quarks and mesons serve as fundamental degrees of freedom, and the confinement of quarks by an effective confining potential results in the formation of baryons.
+The constituent quarks and the mesons obtain their masses through the inclusion of properties such as spontaneous and explicit (for pseudoscalar mesons) symmetry breaking fortyfour .
 In the CQMF model, the effective Lagrangian density accounting for various interaction terms is given as fortyfour ; fortyfive ; fiftythree
-${\cal L}_{{\rm eff}}\,=\,{\cal L}_{q0}\,+\,{\cal L}_{qm}\,+\,{\cal L}_{VV}\,+\,{\cal L}_{\Sigma\Sigma}\,+\,{\cal L}_{\chi SB}\,+\,{\cal L}_{c},$ (1)
-where $\,{\cal L}_{q0}=\bar{q}i\gamma^{\mu}\partial_{\mu}q\,$ is the kinetic term for the free part of massless quarks and ${\cal L}_{qm}\,$ is the quark-meson interaction term that remains invariant under chiral $SU(3)$ transformations.
-${\cal L}_{qm}\,$ can be written as fortyfour ; fortysix ; fiftythree ; fiftyfour
-$\mathcal{L}_{qm}$ $=g_{\mathrm{s}}\left(\bar{\Psi}_{L}M\Psi_{R}+\bar{\Psi}_{R}M^{\dagger}\Psi_{L}\right)-g_{v}\left(\bar{\Psi}_{L}\gamma^{\mu}l_{\mu}\Psi_{L}+\bar{\Psi}_{R}\gamma^{\mu}r_{\mu}\Psi_{R}\right)$ (2)
-$=\frac{g_{\mathrm{s}}}{\sqrt{2}}\bar{\Psi}\left(\sum_{a=0}^{8}s_{a}\lambda_{a}+\mathrm{i}\gamma^{5}\sum_{a=0}^{8}p_{a}\lambda_{a}\right)\Psi-\frac{g_{v}}{2\sqrt{2}}\bar{\Psi}\left(\gamma^{\mu}\sum_{a=0}^{8}v_{\mu}^{a}\lambda_{a}-\gamma^{\mu}\gamma^{5}\sum_{a=0}^{8}a_{\mu}^{a}\lambda_{a}\right)\Psi.$
+${\cal L}_{{\rm effective}}\,=\,{\cal L}_{q0}\,+\,{\cal L}_{qm}\,+\,{\cal L}_{VV}\,+\,{\cal L}_{\Sigma\Sigma}\,+\,{\cal L}_{\chi SB}\,+\,{\cal L}_{cp},$ (1)
+where $\,{\cal L}_{q0}=i\bar{q}\gamma^{\mu}\partial_{\mu}q\,$ represents the kinetic term for quarks having zero mass, while ${\cal L}_{qm}\,$ accounts for the interaction between quarks and mesons, which remains invariant when undergoing chiral $SU(3)$ transformations.
+${\cal L}_{qm}\,$ is expressed as fortyfour ; fortysix ; fiftythree ; fiftyfour
+$\mathcal{L}_{qm}$ $=g_{\mathrm{s}}\left(\bar{\Psi}_{R}{M_{n}}^{\dagger}\Psi_{L}+\bar{\Psi}_{L}{M_{n}}\Psi_{R}\right)-g_{v}\left(\bar{\Psi}_{L}\gamma^{\mu}L_{\mu}\Psi_{L}+\bar{\Psi}_{R}\gamma^{\mu}r_{\mu}\Psi_{R}\right)$ (2)
+$=\frac{g_{\mathrm{s}}}{\sqrt{2}}\bar{\Psi}\left(\mathrm{i}\gamma^{5}\sum_{b=0}^{8}p_{b}\lambda_{b}+\sum_{b=0}^{8}s_{b}\lambda_{b}\right)\Psi-\frac{g_{v}}{2\sqrt{2}}\bar{\Psi}\left(\gamma^{\mu}\sum_{b=0}^{8}v_{\mu}^{b}\lambda_{b}-\gamma^{\mu}\gamma^{5}\sum_{b=0}^{8}a_{\mu}^{b}\lambda_{b}\right)\Psi.$
 In eq.
-2, $\Psi=$ $\begin{pmatrix}u\\ d\\ s\end{pmatrix},$ and $g_{\mathrm{s}}$ and $g_{v}$ are the respective scalar and vector coupling constants.
-The relations between various quark-meson coupling constants are given as fortysix
+2, $g_{\mathrm{s}}$ and $g_{v}$ represent scalar and vector coupling constants, respectively and $\Psi=$ $\begin{pmatrix}u\\ d\\ s\end{pmatrix}$.
+The relationships between these coupling constants are given as fortysix
 $\frac{g_{s}}{\sqrt{2}}=$ $g_{\delta}^{u}=-g_{\delta}^{d}=g_{\sigma}^{u}=g_{\sigma}^{d}=\frac{1}{\sqrt{2}}g_{\zeta}^{s},~{}~{}~{}~{}~{}g_{\delta}^{s}=g_{\sigma}^{s}=g_{\zeta}^{u}=g_{\zeta}^{d}=0\,,$ (3)
-$\frac{g_{v}}{2\sqrt{2}}=g_{\rho}^{u}=-g_{\rho}^{d}=g_{\omega}^{u}=g_{\omega}^{d},~{}~{}~{}~{}~{}~{}~{}g_{\omega}^{s}=g_{\rho}^{s}=0.$ (4)
-It is important to note that spin-0 scalar ($\Sigma$) and pseudoscalar ($\Pi$) meson nonets can be expressed as fortyfour ; fortysix
-$M\left(M^{\dagger}\right)=\Sigma\pm i\Pi=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}\left(s^{a}\pm i\pi^{a}\right)\lambda^{a},$ (5)
-where $s^{a}$ and $\pi^{a}$ represent the nonets of scalar and pseudoscalar mesons, respectively, while $\lambda^{a}$ are Gell-Mann matrices with $\lambda^{0}=\sqrt{\frac{2}{3}}I$.
-The spin-1 mesons can also be defined in a similar way as
-$l_{\mu}\left(r_{\mu}\right)=\frac{1}{2}\left(V_{\mu}\pm A_{\mu}\right)=\frac{1}{2\sqrt{2}}\sum_{a=0}^{8}\left(v_{\mu}^{a}\pm a_{\mu}^{a}\right)\lambda^{a}.$ (6)
-In the above expression, $v_{\mu}^{a}$ and $a_{\mu}^{a}$ are nonets of vector and pseudovector mesons, respectively. The physical states of scalar and vector meson nonets are given by
-$\Sigma=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}s_{a}\lambda^{a}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\sigma+\delta^{0}\right)&\delta^{+}&\kappa^{*+}\\ \delta^{-}&\frac{1}{\sqrt{2}}\left(\sigma-\delta^{0}\right)&\kappa^{*0}\\ \kappa^{*-}&\bar{\kappa}^{*0}&\zeta\end{array}\right),$ (10)
+$\frac{g_{v}}{2\sqrt{2}}=$ $g_{\rho}^{u}=-g_{\rho}^{d}=g_{\omega}^{u}=g_{\omega}^{d}=\frac{g_{\phi}^{s}}{\sqrt{2}},~{}~{}~{}~{}~{}~{}~{}g_{\omega}^{s}=g_{\rho}^{s}=g_{\phi}^{u}=g_{\phi}^{d}=0.$ (4)
+It is important to note that spin-0 scalar, $\Sigma$ and pseudo-scalar meson nonets $\Pi$, are expressed as fortyfour ; fortysix
+$M_{n}\left({M_{n}}^{\dagger}\right)=\Sigma\pm i\Pi=\frac{1}{\sqrt{2}}\sum_{b=0}^{8}\left(s^{b}\pm i\pi^{b}\right)\lambda^{b},$ (5)
+where $s^{b}$ and $\pi^{b}$ represent scalar and pseudo-scalar meson nonents, respectively, while $\lambda^{b}$ represents Gell-Mann matrices with $\lambda^{0}=\sqrt{\frac{2}{3}}I$.
+The spin-1 mesons can also be defined by the relation
+$L_{\mu}\left(r_{\mu}\right)=\frac{1}{2}\left(V_{\mu}\pm A_{\mu}\right)=\frac{1}{2\sqrt{2}}\sum_{b=0}^{8}\left(v_{\mu}^{b}\pm a_{\mu}^{b}\right)\lambda^{b}.$ (6)
+In the above expression, $v_{\mu}^{b}$ and $a_{\mu}^{b}$ are vector and pseudovector meson nonents, respectively. The expressions representing scalar and vector meson nonets are
+$\Sigma=\frac{1}{\sqrt{2}}\sum_{b=0}^{8}s_{b}\lambda^{b}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\delta^{0}+\sigma\right)&\delta^{+}&\kappa^{*+}\\ \delta^{-}&\frac{1}{\sqrt{2}}\left(\sigma-\delta^{0}\right)&\kappa^{*0}\\ \kappa^{*-}&\bar{\kappa}^{*0}&\zeta\end{array}\right),$ (10)
 and
-$V_{\mu}=\frac{1}{\sqrt{2}}\sum_{a=0}^{8}v_{\mu}^{a}\lambda^{a}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\omega_{\mu}+\rho_{\mu}^{0}\right)&\rho_{\mu}^{+}&K_{\mu}^{*+}\\ \rho_{\mu}^{-}&\frac{1}{\sqrt{2}}\left(\omega_{\mu}-\rho_{\mu}^{0}\right)&K_{\mu}^{*0}\\ K_{\mu}^{*-}&\bar{K}_{\mu}^{*0}&\phi_{\mu}\end{array}\right),$ (14)
+$V_{\mu}=\frac{1}{\sqrt{2}}\sum_{b=0}^{8}v_{\mu}^{b}\lambda^{b}=\left(\begin{array}[]{lcc}\frac{1}{\sqrt{2}}\left(\omega_{\mu}+\rho_{\mu}^{0}\right)&\rho_{\mu}^{+}&K_{\mu}^{*+}\\ \rho_{\mu}^{-}&\frac{1}{\sqrt{2}}\left(\omega_{\mu}-\rho_{\mu}^{0}\right)&K_{\mu}^{*0}\\ K_{\mu}^{*-}&\bar{K}_{\mu}^{*0}&\phi_{\mu}\end{array}\right),$ (14)
 respectively.
-The term, ${\cal L}_{VV}$ in eq.
-1 represents the chiral-invariant vector meson self interaction terms.
-Within the framework of mean field approximation ${\cal L}_{VV}$ is given by fortyfour ; fiftythree ; fiftyfour
+The chiral invariant term ${\cal L}_{VV}$ in eq. 1, incorporates the mass term for the vector mesons as well as higher order self interactions and is written as fortyfour
+${\cal L}_{VV}=\frac{1}{2}m_{V}^{2}\frac{\chi^{2}}{\chi_{0}^{2}}\operatorname{Tr}V_{\mu}V^{\mu}+2g_{4}\operatorname{Tr}\left(V_{\mu}V^{\mu}\right)^{2}.$ (15)
+The expression for ${\cal L}_{VV}$ under mean field approximation comes out to be fortyfour ; fiftythree ; fiftyfour
 ${\cal L}_{VV}=\frac{1}{2}\,\frac{\chi^{2}}{\chi_{0}^{2}}\left(m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2}+m_{\phi}^{2}\phi^{2}\right)+g_{4}\left(\omega^{4}+6\omega^{2}\rho^{2}+\rho^{4}+2\phi^{4}\right).$ (16)
 For the scalar mesons, their self interaction terms are written as (fourth term in eq. 1)
-${\cal L}_{\Sigma\Sigma}=$ $-\frac{1}{2}\,k_{0}\chi^{2}\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)+k_{1}\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)^{2}+k_{2}\left(\frac{\sigma^{4}}{2}+\frac{\delta^{4}}{2}+3\sigma^{2}\delta^{2}+\zeta^{4}\right)$
+${\cal L}_{\Sigma\Sigma}=$ $-\frac{1}{2}\,k_{0}\chi^{2}\left(\delta^{2}+\zeta^{2}+\sigma^{2}\right)+k_{1}\left(\delta^{2}+\zeta^{2}+\sigma^{2}\right)^{2}+k_{2}\left(3\sigma^{2}\delta^{2}+\frac{\delta^{4}}{2}+\frac{\sigma^{4}}{2}+\zeta^{4}\right)$
 $+k_{3}\chi\left(\sigma^{2}-\delta^{2}\right)\zeta-k_{4}\chi^{4}-\frac{1}{4}\chi^{4}{\rm ln}\frac{\chi^{4}}{\chi_{0}^{4}}+\frac{\xi}{3}\chi^{4}{\rm ln}\left(\left(\frac{\left(\sigma^{2}-\delta^{2}\right)\zeta}{\sigma_{0}^{2}\zeta_{0}}\right)\left(\frac{\chi^{3}}{\chi_{0}^{3}}\right)\right).$ (17)
-The parameters $k_{0},k_{1},k_{2},k_{3}$ and $k_{4}$ appearing in the above Lagrangian density are calculated using masses of $\pi$ and $K$ as well as the average mass of $\eta$ and $\eta^{{}^{\prime}}$ mesons.
-The vacuum expectation values of scalar meson fields $\sigma$ and $\zeta$, i.e.
-$\sigma_{0}$ and $\zeta_{0}$, can be expressed in terms of pion and kaon leptonic decay constants $f_{\pi}$ and $f_{K}$, through the relations
+The parameters $k_{i=0,1,2,3,4}$ appearing in the above Lagrangian density are calculated using masses of $\pi$ and $K$ as well as the average of the $\eta$ and $\eta^{{}^{\prime}}$ meson masses.
+The expectation values of scalar meson fields $\sigma$ and $\zeta$, in free space i.e.
+$\sigma_{0}$ and $\zeta_{0}$, can be expressed with pion and kaon leptonic decay constants $f_{\pi}$ and $f_{K}$, with the expressions
 $\sigma_{0}=-f_{\pi}~{}~{}{\rm and}~{}~{}~{}~{}\zeta_{0}=\frac{1}{\sqrt{2}}\left(f_{\pi}-2f_{K}\right),$ (18)
 where $f_{\pi}=92.8$ MeV and $f_{K}=115$ MeV.
-The vacuum value of dilaton field $\chi_{0}=254.6$ MeV and the coupling constant $g_{4}=37.4$, obtained from fit reasonable effective nucleon mass fortyfour ; fiftythree ; fiftyfour .
-The last three terms of eq.
-16 are introduced in the model to incorporate the trace anomaly property of QCD, which leads to the trace of energy momentum tensor being proportional to fourth power of dilaton field $\chi$ fortynine .
-The parameter $\xi$, originating from the logarithmic term used in eq.
-16, is usually determined from QCD $\beta$-function at one loop level for three colors and flavors fortynine .
-The fifth term of the effective Lagrangian density of eq.
-1 is the explicit chiral symmetry breaking term, ${\cal L}_{\chi SB}$, and is expressed as
+In absence of nuclear matter, the dilaton field $\chi_{0}=254.6$ MeV and the coupling constant $g_{4}=37.4$, have been obtained from a reasonable fit of effective nucleon mass fortyfour ; fiftythree ; fiftyfour .
+The concluding three terms of eq.
+17 incorporates the trace anomaly property of QCD, which leads to the trace of energy momentum tensor being proportional to the dilaton field $\chi$ raised to the power of 4 fortynine .
+The explicit chiral symmetry breaking term ${\cal L}_{\chi SB}$ in eq.
+1 is expressed through the relation
 ${\cal L}_{\chi SB}=\frac{\chi^{2}}{\chi_{0}^{2}}\left[m_{\pi}^{2}f_{\pi}\sigma+\left(\sqrt{2}\,m_{K}^{2}f_{K}-\frac{m_{\pi}^{2}}{\sqrt{2}}f_{\pi}\right)\zeta\right].$ (19)
-The above expression is introduced in chiral effective models to account for the non-vanishing pseudoscalar meson masses and it satisfies the partial conserved axial-vector current relations for $\pi$ and $K$ mesons fortyfour ; fortysix ; fortyseven .
-The confinement of constituent quarks inside the baryons, due to the presence of a confining scalar-vector potential $\chi_{c}$, is introduced through the last term in the Lagrangian density, in eq.
-1, and is written as
-${\cal L}_{c}=-\bar{\psi}\chi_{c}\psi,$ (19)
-where the confining potential $\chi_{c}$ is given by
+The above expression accounts for the non-zero masses of pseudoscalar mesons and agrees well with the partial conserved axial-vector current relations of $\pi$ and $K$ mesons fortyfour ; fortysix ; fortyseven .
+Unlike the models which uses the concept of spherical-bag boundary, like MIT bag Model, the CQMF model treats the constituent quarks to be moving inside the baryons forming a bounded system, by virtue of a confining potential $\chi_{c}$.
+The use of an average confining potential with quarks moving around the baryon core, has been fruitful in improving the static baryonic properties like magnetic moment of proton Barik1985 .
+The last term in the Lagrangian density, in eq.
+1, incorporates this as
+${\cal L}_{cp}=-\bar{\psi}\chi_{c}\psi,$ (20)
+and the confining potential $\chi_{c}$, is expressed as
 $\chi_{c}(r)=\frac{1}{4}k_{c}\,r^{2}(1+\gamma^{0})\,.$ (21)
-In eq. 20 above the coupling constant $k_{c}$ is taken to be $98\,\text{MeV}.\text{fm}^{-2}$.
-In order to investigate the properties of asymmetric strange matter at finite temperature and density, we are using mean field approximation fortyfour .
-Under the influence of meson mean field, we express the Dirac equation for the quark field $\Psi_{qi}$ as
-$\left[-i\vec{\alpha}\cdot\vec{\nabla}+\chi_{c}(r)+\beta m_{q}^{*}\right]\Psi_{qi}=e_{q}^{*}\Psi_{qi}.$ (21)
-Here the quark $q$ ($q=u,d,s$) is in a baryon of type $i$ ($i=N,\Lambda,\Sigma,\Xi$)  and $\vec{\alpha}$ , $\beta$  are the usual Dirac matrices. The effective quark mass $m_{q}^{*}$ is defined as
-$m_{q}^{*}=-g_{\sigma}^{q}\sigma-g_{\zeta}^{q}\zeta-g_{\delta}^{q}I^{3q}\delta+m_{q0},$ (22)
-At zero baryon density, the masses of constituent quarks $u$, $d$, $s$ are expressed in terms of vacuum expectation values $\sigma_{0}$ and $\zeta_{0}$ of meson scalar fields $\sigma$ and $\zeta$, respectively.
-For the light $u$ and $d$ quarks, we have
-$m_{u}=m_{d}=-g_{\sigma}^{q}\sigma_{0}=-\frac{g_{s}}{\sqrt{2}}\sigma_{0}.$ (23)
-The values of coupling constant $g_{s}$ and confining parameter $k_{c}$ are fitted to obtain binding energy of $-16.0$ MeV at nuclear saturation density, $\rho_{0}=0.16$ $\text{fm}^{-3}$.
+In eq. 21 above the coupling constant $k_{c}$ is considered to be $98\,\text{MeV}.\text{fm}^{-2}$.
+In light of mean field approximation, we are to investigate the asymmetric strange matter properties at finite temperature and density, we are using fortyfour .
+In the presence of meson mean field, we express the Dirac equation for the quark field $\Psi_{qi}$ as
+$\left[-i\vec{\alpha}\cdot\vec{\nabla}+\beta m_{q}^{*}+\chi_{c}(r)\right]\Psi_{qi}=e_{q}^{*}\Psi_{qi}.$ (22)
+Here the quark $q$ is in a baryon of type $i$ and $\vec{\alpha}$ , $\beta$  are the Dirac matrices. By definition, the in-medium constituent quark mass $m_{q}^{*}$ is
+$m_{q}^{*}=-g_{\sigma}^{q}\sigma-g_{\zeta}^{q}\zeta-g_{\delta}^{q}I^{3q}\delta+m_{q0}.$ (23)
+In free space, the masses of constituent quarks $u$, $d$, $s$ are expressed in terms of $\sigma_{0}$ and $\zeta_{0}$, by the relation
+$m_{u}=m_{d}=-\frac{g_{s}}{\sqrt{2}}\sigma_{0}=-g_{\sigma}^{q}\sigma_{0}.$ (24)
+The coupling constant $g_{s}$ and confining parameter $k_{c}$ are fitted to obtain binding energy of $-16.0$ MeV at nuclear saturation density, $\rho_{0}=0.16$ $\text{fm}^{-3}$ fortyfive .
 This fit provides the value of effective mass of light quarks to be equal to $256$ MeV.
-Here $m_{q0}$ is zero for non-strange $u$ and $d$ quarks, whereas for strange $s$ quark $m_{q0}=\Delta m=77$ MeV.
-The term $\Delta m$ is introduced through the Lagrangian density ${\cal L}_{\Delta m}=-\Delta m\bar{\psi}m_{1}\psi$, where $S_{1}=\frac{1}{3}\left(1-\lambda_{8}\sqrt{3}\right)$.
+Here $m_{q0}$ is zero for non-strange $u$ and $d$ quarks, while for $s$ quark $m_{q0}=\Delta m=77$ MeV.
 In eq.
-21, $e_{q}^{*}$ in the Dirac equation is the effective energy of quarks and is expressed as fortyfour
+22, $e_{q}^{*}$ in the Dirac equation is the effective energy of quarks and is expressed as fortyfour
 $e_{q}^{*}=e_{q}-g_{\omega}^{q}\omega-g_{\rho}^{q}I^{3q}\rho-g_{\phi}^{q}\phi.$ (25)
-The in-medium mass of a baryon is expressed in terms of its effective energy $E^{*}_{i}$ and spurious center of mass momentum $p_{icm}$ as in Barik1985 ; Barik2013 ,
-$M^{*}_{i}=\sqrt{E^{*2}_{i}-<p^{*2}_{icm}>}.$ (25)
+The in-medium mass of a baryon is expressed as,
+$M^{*}_{i}=\sqrt{E^{*2}_{i}-<p^{*2}_{icm}>},$ (26)
+in terms of its in-medium energy $E^{*}_{i}$ and spurious center mass motion $p_{icm}$ Barik1985 ; Barik2013 .
+Spuirous center of mass momentum $p_{icm}$ corrections are added in expression for $M_{i}^{*}$ to further improve the overlooked independent quark motion inside the baryons.
+The explicit expression for $<p^{*2}_{icm}>$ being summation over average value of square of each constituent quark momentum ($<p^{*2}_{cm}>_{q}$) Barik1985 , i.e.,
+$<p^{*2}_{icm}>=\sum_{q}<p^{*2}_{cm}>_{q}=\sum_{q}\frac{11e_{q}^{*}+m_{q}^{*}}{6(3e_{q}^{*}+m_{q}^{*})}(e_{q}^{*2}-m_{q}^{*2}).$ (27)
+As can be seen from the above equation, spurious center of mass motion corrections are density and temperature dependent.
+In Ref.
+fortyseven , spurious center of mass motion corrections are considered as independent of medium effects.
 Further the in-medium energies of $i^{th}$ baryon can be related to the in-medium energies of quarks and number of q type constituent quarks $n_{qi}$ in $i^{th}$ baryon by the expression
 $E^{*}_{i}=\sum_{q}n_{qi}e_{q}^{*}+E_{ispin}.$ (28)
-Here, $E_{ispin}$ is a correction term in the effective energy of a baryon that accounts for spin-spin interaction between constituent quarks.
-The $E_{ispin}$ is determined so as to obtain the correct baryon masses at zero density.
+Here, $E_{ispin}$ accounts for spin-spin interaction of constituent quarks and serves as a correction term in the effective energy of a baryon.
+The $E_{ispin}$ is computed so as to obtain the correct baryon masses at zero density.
 Finally, the thermodynamic potential $\Omega$ describing the isospin asymmetric strange hadronic matter at finite temperature and baryon density is written as fortyfour
-$\Omega=-\frac{k_{B}T}{(2\pi)^{3}}\sum_{i}\gamma_{i}\int_{0}^{\infty}d^{3}k\left\{\ln\left(1+e^{-\left[E_{i}^{*}(k)-\nu_{i}^{*}\right]/k_{B}T}\right)+\ln\left(1+e^{-\left[E_{i}^{*}(k)+\nu_{i}^{*}\right]/k_{B}T}\right)\right\}-\mathcal{L}_{M}-\mathcal{V}_{\text{vac }},$ (27)
-where the mesonic Lagrangian is $\mathcal{L}_{M}=\mathcal{L}_{VV}+\mathcal{L}_{\Sigma\Sigma}+\mathcal{L}_{\chi SB}$ and we have $E^{*}_{i}(k)=\sqrt{M_{i}^{*2}+k^{2}}$.
+$\Omega=-\frac{k_{B}T}{(2\pi)^{3}}\sum_{i}\gamma_{i}\int_{0}^{\infty}\left\{\ln\left(e^{-\left[E_{i}^{*}(k)-\nu_{i}^{*}\right]/k_{B}T}+1\right)+\ln\left(e^{-\left[E_{i}^{*}(k)+\nu_{i}^{*}\right]/k_{B}T}+1\right)\right\}d^{3}k-\mathcal{L^{\prime}}-\mathcal{V}_{\text{vac }},$ (29)
+where the mesonic Lagrangian is $\mathcal{L^{\prime}}=\mathcal{L}_{VV}+\mathcal{L}_{\Sigma\Sigma}+\mathcal{L}_{\chi SB}$ and we have $E^{*}_{i}(k)=\sqrt{M_{i}^{*2}+k^{2}}$.
 In the above equation, the degeneracy factor is $\gamma_{i}=2$ for baryons and the summation is over hyperons and nucleons.
 In eq.
-27, the effective chemical potential $\nu_{i}^{*}$ is expressed in terms of the usual baryon chemical potential $\mu_{i}$ as fortyfour ; fiftythree ; fiftyfour ; sixtytwo
+29, the effective chemical potential $\nu_{i}^{*}$ is expressed in terms of the usual baryon chemical potential $\mu_{i}$ as fortyfour ; fiftythree ; fiftyfour ; sixtytwo
 $\nu_{i}^{*}=\mu_{i}-g_{\phi}^{i}\phi-g_{\omega}^{i}\omega-g_{\rho}^{i}I^{3i}\rho.$ (30)
-From eq. 27 the equations of motion of scalar and vector meson fields are calculated by minimizing the effective thermodynamic potential as (see APPENDIX)
+From eq. 29 the equations of motion of scalar and vector meson fields are calculated by minimizing the effective thermodynamic potential as
 $\frac{\partial\Omega}{\partial\sigma}=\frac{\partial\Omega}{\partial\zeta}=\frac{\partial\Omega}{\partial\chi}=\frac{\partial\Omega}{\partial\delta}=\frac{\partial\Omega}{\partial\omega}=\frac{\partial\Omega}{\partial\rho}=\frac{\partial\Omega}{\partial\phi}=0.$ (31)
-Equation 29 is solved for finite values of temperature and density of the medium.
+Equation 31 is solved for finite values of temperature and densities of the medium.
+Note that the confining potential $\chi_{c}$ contribute to the effective energy $e_{q}^{*}$ and hence, masses $m_{q}^{*}$ of quarks through the solution of Dirac equation.
+The effective quark energies are used in calculations of baryon energy which further enter into the coupled system of equations for scalar and vector fields through the definition of scalar and vector density of baryons fortyfour .
 In order to account for isospin asymmetry in the hadronic medium a parameter $i_{a}=-\frac{\Sigma_{i}I_{3i}\rho_{i}}{\rho_{B}}$ is introduced, where $I_{3i}$ represents the third component of isopin and $\rho_{B}$ is baryonic density of the medium fiftyfour .
 To take into account hyperons in the medium, we define strangeness fraction $f_{s}=\frac{\Sigma_{i}\left|s_{i}\right|\rho_{i}}{\rho_{B}}$, here $|s_{i}|$ denotes the number of $s$ quarks in $i^{th}$ baryon.
+Note that in the present manuscript, the difference in the $\frac{1}{2}^{-}$ baryon resonances and positive parity octet baryons comes through the fitting of $E_{ispin}$ value to their respective vacuum masses.
+The calculations of effective masses of $\frac{1}{2}^{-}$ baryon resonances proceed in two steps: first we solve the system of equations corresponding to the scalar ($\sigma,\zeta$ and $\delta$) and vector fields ($\omega,\rho$ and $\phi$) in the strange hadronic medium (considering only positive parity octet baryons and utilizing Eqs.
+(26) and (28) for in-medium effective masses and energies) at finite density and temperature (see Eq.
+(31)).
+This gives us the density and temperature dependent values of scalar and vector fields which are used in the second step to obtain the effective masses of $\frac{1}{2}^{-}$ baryon resonances.
+For this, first we use the vacuum ($\rho_{B}=0,T=0,i_{a}=0,f_{s}=0$) values of fields calculated in first step, in Eqs.
+(26) and (28 and fit $E_{ispin}$ to vacuum masses of $\frac{1}{2}^{-}$ baryon resonances.
+Using these $E_{ispin}$ values and density and temperature dependent values of scalar and vector fields calculated in the first step (strange matter), the in-medium masses of $\frac{1}{2}^{-}$ baryon resonances are calculated using Eqs.
+(26) in the strange hadronic medium.
 We have eq.
 26 which addresses our aim to compute the effective baryon masses, by taking into account the effective constituent quark masses, baryon density and the spurious center of mass motion.
-In the upcoming section we shift the trajectory of our discussion towards the calculation of effective baryonic magnetic moments, by incorporating the medium modified constituent quark masses and the effective baryon masses.
-III Magnetic Moment of ${\frac{1}{2}}^{-}$ baryon resonances
-In the present section, methodology used to calculate the magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances has been discussed.
-The formalism of chiral constituent quark model ($\chi\mathrm{CQM}$) thirtyone , has been used to calculate the magnetic moment of the baryons through the following fluctuation process eight ; fiftyfive
-$q^{\uparrow\downarrow}\rightarrow\mathrm{GB}+q^{\prime\downarrow\uparrow}\rightarrow{\left(q\overline{q}^{\prime}\right)+q^{\prime\downarrow\uparrow}},$ (30)
-where GB represents the emitted Goldstone boson and $q\overline{q}^{\prime}+q^{\prime}$ constitute the sea quarks nineteen ; twenty ; twentyone ; twentytwo ; twentythree .
-The $\chi\mathrm{CQM}$ model incorporates the concept of confinement and chiral symmetry breaking twelve ; fiftysix , where the massless quarks acquire mass through spontaneous breaking of chiral symmetry twelve .
-In this approach we consider the emission of a GB, which further splits into a $q\overline{q}^{\prime}$ pair.
-The coupling between the quark and GBs is described by the effective interaction Lagrangian,
-$\mathcal{L}=g_{8}\overline{\mathbf{q}}\left(\Phi+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}I\right)\mathbf{q}=g_{8}\overline{\mathbf{q}}\left(\Phi^{\prime}\right)\mathbf{q}.$ (31)
-This quark-GB coupling represented in this Lagrangian is weak enough to consider the fluctuation in eq.
-30 as a small perturbation and the higher order fluctuations can therefore be neglected.
-The process of emission of GB from the quark gives rise to the probability of a spin-flip of the quark.
-In eq.
-31, $\zeta^{\prime}=g_{1}/g_{8},$ where $g_{1}$ and $g_{8}$ denote the coupling constants for the singlet and octet GBs, respectively and $I$ is a $3\times 3$ identity matrix.
-The matrix of the GBs can be written as
+In the upcoming section we shift the trajectory of our discussion towards the calculation of effective baryonic magnetic moments, by incorporating the effective in-medium masses of constituent quarks and the baryons under study.
+2.2 Magnetic Moment of ${\frac{1}{2}}^{-}$ baryon resonances
+The formalism of chiral constituent quark model ($\chi\mathrm{CQM}$) thirtyone , has been implemented to calculate the magnetic moment of the baryons through the following fluctuation process eight ; fiftyfive
+$q^{\uparrow\downarrow}\rightarrow\mathrm{NGB}+q^{\prime\downarrow\uparrow}\rightarrow{\left(q\overline{q}^{\prime}\right)+q^{\prime\downarrow\uparrow}},$ (32)
+where the emitted Nambu-Goldstone boson (NGB) furthermore splits into a $q\overline{q}^{\prime}$ pair, and $q\overline{q}^{\prime}+q^{\prime}$ constitute the sea quarks nineteen ; twenty ; twentyone ; twentytwo ; twentythree .
+In $\chi\mathrm{CQM}$ model the massless confined quarks obtain its mass by virtue of spontaneous chiral symmetry breaking twelve ; fiftysix .
+The coupling between the quark and NGBs is described by the effective interaction Lagrangian
+$\mathcal{L}=g_{8}\overline{\mathbf{q}}\left(\Phi^{\prime}\right)\mathbf{q}=g_{8}\overline{\mathbf{q}}\left(\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}I+\Phi\right)\mathbf{q},$ (33)
+where $\zeta^{\prime}=g_{1}/g_{8},$ where $g_{1}$ and $g_{8}$ denote the coupling constants for the singlet and octet NGBs respectively and $I$ is an order 3 identity matrix.
+This quark-NGB coupling represented in this Lagrangian is quite feeble and hence, the fluctuation in eq.
+32 is practically taken to be a minor perturbation.
+The spin-flip process is an outcome of the emission of NGB.
+The NGB matrix is defined by
 $\Phi^{\prime}=\left(\begin{array}[]{ccc}\frac{\pi^{0}}{\sqrt{2}}+\beta\frac{\eta}{\sqrt{6}}+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}&\pi^{+}&\alpha K^{+}\\ \pi^{-}&-\frac{\pi^{0}}{\sqrt{2}}+\beta\frac{\eta}{\sqrt{6}}+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}&\alpha K^{0}\\ \alpha K^{-}&\alpha\bar{K}^{0}&-\beta\frac{2\eta}{\sqrt{6}}+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}\end{array}\right)$ (34)
@@
 $q=\left(\begin{array}[]{l}u\\ d\\ s\end{array}\right).$
-The $SU(3)$ chiral symmetry breaking is introduced in $\chi\mathrm{CQM}$ by considering that $m_{s}>m_{u,d}$ and the masses of GBs as non-degenerate $\left(m_{\eta^{\prime}}>m_{K,\eta}>m_{\pi}\right)$ fiftyseven .
-The parameter $a=\left|g_{8}\right|^{2}$ is denoted as the transition probability of chiral fluctuation $u(d)\rightarrow$ $d(u)+\pi^{+(-)}$, whereas $a\alpha^{2},a\beta^{2}$ and $a\zeta^{\prime 2}$ denote the probabilities of transitions of $u(d)\rightarrow s+K^{-(0)},u(d,s)\rightarrow$ $u(d,s)+\eta$ and $u(d,s)\rightarrow u(d,s)+\eta^{\prime}$, respectively nineteen ; twenty ; twentyone ; twentytwo ; twentythree .
-By thoroughly analysing the spin and flavor distribution functions of proton, the best fitted set of parameters crucial to our calculations are nineteen ; twentythree ; twentyfive ; twentyeight ; twentynine ; thirty ; thirtyone : $a=0.12,\alpha=\beta=0.45,\eta=-0.15$.
+The $SU(3)$ chiral symmetry breaking is introduced in $\chi\mathrm{CQM}$ by treating the masses of NGBs as non-degenerate and considering that $m_{s}>m_{u,d}$ fiftyseven .
+The parameter for the transition probability of chiral fluctuations: $u(d)\rightarrow$ $d(u)+\pi^{+(-)}$ is denoted by $a(=\left|g_{8}\right|^{2})$, $u(d)\rightarrow s+K^{-(0)}$ by $a\alpha^{2}$, $u(d,s)\rightarrow$ $u(d,s)+\eta$ by $a\beta^{2}$ and $u(d,s)\rightarrow u(d,s)+\eta^{\prime}$ is represented by $a\zeta^{\prime 2}$ nineteen ; twenty ; twentyone ; twentytwo ; twentythree .
+By thoroughly analyzing the spin and flavor distribution functions of proton, the best fitted set of parameters crucial to our calculations are nineteen ; twentythree ; twentyfive ; twentyeight ; twentynine ; thirty ; thirtyone : $a=0.12,\alpha=\beta=0.45,\zeta^{\prime}=-0.15$.
 The effective baryonic magnetic moment $\mu^{*}_{B}$ is a summation of the contributions due to spin polarization $\mu^{*S}_{B}$ and the angular momentum polarization $\mu^{*L}_{B}$, expressed as fiftythree ; fiftyfour
 $\mu^{*}_{B}=\mu^{*S}_{B}+\mu^{*L}_{B}.$ (35)
-Supposedly B ¿ is a SU(6) spin-flavor wave function of a baryon, then we have
-$\left\langle B\left|\mu^{*S}_{B}\right|B\right\rangle=\Delta u\mu^{*}_{u}+\Delta d\mu^{*}_{d}+\Delta s\mu^{*}_{s},$ (34)
-$\left\langle B\left|\mu^{*L}_{B}\right|B\right\rangle=\Delta u^{(1)}\mu^{*}_{u}+\Delta d^{(1)}\mu^{*}_{d}+\Delta s^{(1)}\mu^{*}_{s}.$ (35)
-Here, $\mu^{*}_{q}(q=u,d,s)$ is the effective quark magnetic moment.
-The spin and orbital angular momentum polarizations are quantitatively comprehended from $\Delta q=q^{\uparrow}-q^{\downarrow}$ and $\Delta q^{(1)}=q^{(1)}-q^{(-1)}$, respectively.
-The polarizations can be calculated using the number operator $\mathcal{N}$ (=$\mathcal{N}^{\uparrow\downarrow}+\mathcal{N}^{\pm 1}$) as
-$\widehat{B}\equiv\langle B|\mathcal{N}|B\rangle.$ (36)
-Here $\mathcal{N}^{\uparrow\downarrow}$ and $\mathcal{N}^{(\pm 1)}$ of the number operator $\mathcal{N}$ are expressed as
-$\mathcal{N}^{\uparrow\downarrow}=$ $n_{u\uparrow}u^{\uparrow}+n_{u\downarrow}u^{\downarrow}+n_{d\uparrow}d^{\uparrow}+n_{d^{\downarrow}}d^{\downarrow}+n_{s^{\uparrow}}s^{\uparrow}+n_{s\downarrow}s^{\downarrow},$ (37)
-$\mathcal{N}^{(\pm 1)}=$ $n_{u^{(1)}}u^{(1)}+n_{u(-1)}u^{(-1)}+n_{d^{(1)}}d^{(1)}+n_{d^{(-1)}}d^{(-1)}+n_{s^{(1)}}s^{(1)}+n_{s^{(-1)}}s^{(-1)},$ (38)
-with $n_{q^{\uparrow}}\left(n_{q^{\downarrow}}\right)$ denoting the number of spin up (down) quarks and $n_{q^{(1)}}\left(n_{q^{(-1)}}\right)$ being the number of quarks with the orbital angular momentum projection $m_{L}=1$ $\left(m_{L}=-1\right)$.
-Following the Cheng Li mechanism ten , the calculation of total baryonic magnetic moment involves the contribution of valence and sea quark spin polarizations and orbital angular momentum of sea quarks.
+Following the Cheng Li mechanism ten , the calculation of total baryonic magnetic moment involves the valence and sea quark contributions and the orbital angular momentum of sea quarks.
 The spin part in eq.
-33, can be obtained from contributions due to valence, sea and orbital angular momentum of sea quarks and is given by
+35, comprises of these individual contributions and is given by
 $\mu^{*S}_{B}=\mu^{*S}_{val}+\mu^{*S}_{sea}+\mu^{*S}_{orbit}.$ (36)
@@
 $q^{\uparrow(\downarrow)}\rightarrow-P_{[q,GB]q^{\uparrow(\downarrow)}}+\left|\psi\left(q^{\uparrow(\downarrow)}\right)\right|^{2}.$ (37)
-In eq. 40, $P_{[q,GB]}$ is the probability of emission of GBs from a quark $q^{\uparrow(\downarrow)}$ fiftyeight given by
-$P(u\rightarrow\mathrm{GB})=P(d\rightarrow\mathrm{GB})=\frac{a}{6}\left(9+6\alpha^{2}+\beta^{2}+2\zeta^{\prime 2}\right),$ (41)
-$P(s\rightarrow\mathrm{GB})=\frac{a}{3}\left(6\alpha^{2}+2\beta^{2}+\zeta^{\prime 2}\right),$
-and $\left|\psi\left(q^{\uparrow(\downarrow)}\right)\right|^{2}$ represents the probability of transforming a $q^{\uparrow(\downarrow)}$ quark and it is given by nineteen ; twenty ; twentyone ; twentytwo ; twentythree ; fiftythree ; fiftyfour
-$\left|\psi\left(u^{\uparrow(\downarrow)}\right)\right|^{2}=\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)u^{\uparrow(\downarrow)}+ad^{\uparrow(\downarrow)}+a\alpha^{2}s^{\uparrow(\downarrow)},$ (42)
-$\left|\psi\left(d^{\uparrow(\downarrow)}\right)\right|^{2}=au^{\uparrow(\downarrow)}+\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)d^{\uparrow(\downarrow)}+a\alpha^{2}s^{\uparrow(\downarrow)},$
-$\left|\psi\left(s^{\uparrow(\downarrow)}\right)\right|^{2}=a\alpha^{2}u^{\uparrow(\downarrow)}+a\alpha^{2}d^{\uparrow(\downarrow)}+\frac{a}{3}\left(2\beta^{2}+\zeta^{\prime 2}\right)s^{\uparrow(\downarrow)}.$
-The orbital angular momentum contribution due to sea quarks, is expressed as ten ; fourteen ; twentyeight ; twentynine ; thirty ; thirtyone
-$\mu^{*}\left(q^{\uparrow}\rightarrow q^{\prime\downarrow}\right)=\frac{e_{q}}{2m^{*}_{q}}\left\langle l_{q}\right\rangle+\frac{e_{q}-e_{q\prime}}{2M_{\mathrm{GB}}}\left\langle l_{\mathrm{GB}}\right\rangle,$ (43)
-where $l_{q}\right>=\frac{M_{\mathrm{GB}}}{M_{\mathrm{GB}}+m^{*}_{q}}$ and $l_{\mathrm{GB}}\right>=\frac{m^{*}_{q}}{M_{\mathrm{GB}}+m^{*}_{q}}$.
-The quantities $\left(l_{q},l_{\mathrm{GB}}\right)$ and $\left(m^{*}_{q},M_{\mathrm{GB}}\right)$ stands for the orbital angular momenta and masses of quarks and GBs, respectively.
-The orbital moment of each process in eq.
-43, when multiplied by the probability of occurrence of the process, produces the magnetic moment due to all possible transitions for a given valence quark.
-This can be mathematically expressed as,
-${\left[\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\left(\frac{1}{2}+\frac{\beta^{2}}{6}+\frac{\zeta^{\prime 2}}{3}\right)\mu^{*}\left(u_{\uparrow}\rightarrow u_{\downarrow}\right)\right.$ (44)
-$\left.+\mu^{*}\left(u_{\uparrow}\rightarrow d_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(u_{\uparrow}\rightarrow s_{\downarrow}\right)\right],$
-${\left[\mu^{*}\left(d_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\mu^{*}\left(d_{\uparrow}\rightarrow u_{\downarrow}\right)+\left(\frac{1}{2}+\frac{\beta^{2}}{6}+\frac{\zeta^{\prime 2}}{3}\right)\right.$
-×$\left.\times\mu^{*}\left(d_{\uparrow}\rightarrow d_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(d_{\uparrow}\rightarrow s_{\downarrow}\right)\right],$
-${\left[\mu^{*}\left(s_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\alpha^{2}\mu^{*}\left(s_{\uparrow}\rightarrow u_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(s_{\uparrow}\rightarrow d_{\downarrow}\right)\right.$
-$\left.+\left(\frac{2}{3}\beta^{2}+\frac{\zeta^{\prime 2}}{3}\right)\mu^{*}\left(s_{\uparrow}\rightarrow s_{\downarrow}\right)\right].$
-The orbital moments of the $u,d$ and $s$ quarks can also be expressed in terms of the $\chi\mathrm{CQM}$ parameters $(a,\alpha,\beta,\zeta^{\prime})$, masses of GBs $\left(M_{\pi},M_{\eta\prime},M_{K},M_{\eta}\right)$ and effective quark masses $\left(m^{*}_{u},m^{*}_{d},m^{*}_{s}\right)$, respectively as thirtyone ; fiftythree ; fiftyfour :
-${\left[\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\frac{3m_{u}^{*2}}{2M_{\pi}\left(m_{u}^{*}+M_{\pi}\right)}-\frac{\alpha^{2}\left(M_{K}^{2}-3m_{u}^{*2}\right)}{2M_{K}\left(m_{u}^{*}+M_{K}\right)}\right.+\frac{\beta^{2}M_{\eta}}{6\left(m_{u}^{*}+M_{\eta}\right)}$
-$\left.+\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{3\left(m_{u}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{u},$
-${\left[\mu^{*}\left(d_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{3\left(M_{\pi}^{2}-2m_{d}^{*2}\right)}{4M_{\pi}\left(m^{*}_{d}+M_{\pi}\right)}-\frac{\alpha^{2}M_{K}}{2\left(m_{d}^{*}+M_{K}\right)}-\frac{\beta^{2}M_{\eta}}{12\left(m_{d}^{*}+M_{\eta}\right)}\right.$ (45)
-$\left.-\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{6\left(m_{d}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{d},$
-${\left[\mu^{*}\left(s_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{\alpha^{2}\left(M_{K}^{2}-3m_{s}^{*2}\right)}{2M_{K}\left(m_{s}^{*}+M_{K}\right)}-\frac{\beta^{2}M_{\eta}}{3\left(m_{s}^{*}+M_{\eta}\right)}-\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{6\left(m_{s}^{*}+M_{\eta^{\prime}}\right)}\right]\mu_{s}^{*}.$
-In the present study, we have included the effect of relativistic quark confinement and have calculated the mass adjusted effective magnetic moments of its constituent quarks by the expression fiftyfive
-$\mu_{d}^{*}=-\left(1-\frac{\Delta M}{M_{B}^{*}}\right),\quad\mu_{s}^{*}=-\frac{m_{u}^{*}}{m_{s}^{*}}\left(1-\frac{\Delta M}{M_{B}^{*}}\right),\quad\mu_{u}^{*}=-2\mu_{d}^{*}.$ (46)
-In the above relations, $\Delta M=M_{B,vac}-M_{B}^{*}$ where $M_{B,vac}$ is the vacuum mass of the given baryon and $M_{B}^{*}$ denotes the in-medium mass of this baryon.
-In light of this formalism, we can obtain the spin contribution ($\mu_{B}^{*S}$) towards the baryonic magnetic moment ($\mu^{*}_{B}$).
-The expressions for spin contribution of n* resonance, of $\frac{1}{2}^{-}$ baryon octet, is obtained as in thirtyone
-$\mu_{\text{val}}^{*S}\left(n^{*}\right)=-\left(\frac{2}{9}\mu_{d}^{*}+\frac{1}{9}\mu_{u}^{*}\right)\cos^{2}\theta+\left(\frac{10}{9}\mu_{d}^{*}+\frac{5}{9}\mu_{u}^{*}\right)\sin^{2}\theta\\ -\frac{8}{9}\left(\mu_{d}^{*}-\mu_{u}^{*}\right)\sin\theta\cos\theta,$
-$\hskip 15.6491pt\mu_{\text{sea }}^{*S}\left(n^{*}\right)=$ $\frac{a}{9}\left[\left(5+2\alpha^{2}+\frac{2\beta^{2}}{3}+\frac{4\zeta^{\prime 2}}{3}\right)\mu_{d}^{*}\right.\left.+\left(4+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\mu_{u}^{*}+3\alpha^{2}\mu_{s}^{*}\right]\cos^{2}\theta$
-$-\frac{5a}{9}\left[\left(5+2\alpha^{2}+\frac{2\beta^{2}}{3}+\frac{4\zeta^{\prime 2}}{3}\right)\right.\left.\mu_{d}^{*}+\left(4+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\mu_{u}^{*}+3\alpha^{2}\mu_{s}^{*}\right]\sin^{2}\theta$
-$+\frac{8a}{9}\left(1+\alpha^{2}+\frac{\beta^{2}}{3}+\frac{2\zeta^{\prime 2}}{3}\right)\left(\mu_{d}^{*}-\mu_{u}^{*}\right)\sin\theta\cos\theta,$
-$\mu_{\text{orbit }}^{*S}\left(n^{*}\right)=$ $-\left(\frac{2}{9}\mu^{*}\left(d_{\uparrow}\rightarrow\right)+\frac{1}{9}\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\cos^{2}\theta+\left(\frac{10}{9}\mu^{*}\left(d_{\uparrow}\rightarrow\right)+\frac{5}{9}\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\sin^{2}\theta$ (47)
-$-\frac{8}{9}\left(\mu^{*}\left(d_{\uparrow}\rightarrow\right)-\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right)\sin\theta\cos\theta.$
+In the above expression, the emission probability of NGBs from a quark $q^{\uparrow(\downarrow)}$ is $P_{[q,NGB]}$ which can be fetched from ref.
+fiftyeight , and $\left|\psi\left(q^{\uparrow(\downarrow)}\right)\right|^{2}$ represents the probability of transforming a $q^{\uparrow(\downarrow)}$ quark nineteen ; twenty ; twentyone ; twentytwo ; twentythree ; fiftythree ; fiftyfour .
+The orbital angular momentum contribution due to sea quarks, is given by ten ; fourteen ; twentyeight ; twentynine ; thirty ; thirtyone
+$\mu^{*}\left(q^{\uparrow}\rightarrow q^{\prime\downarrow}\right)=\frac{e_{q}}{2m^{*}_{q}}\left\langle l_{q}\right\rangle+\frac{e_{q}-e_{q\prime}}{2M_{\mathrm{NGB}}}\left\langle l_{\mathrm{NGB}}\right\rangle,$ (38)
+where $l_{q}\right>=\frac{M_{\mathrm{NGB}}}{M_{\mathrm{NGB}}+m^{*}_{q}}$ and $l_{\mathrm{NGB}}\right>=\frac{m^{*}_{q}}{M_{\mathrm{NGB}}+m^{*}_{q}}$.
+The quantities $\left(l_{q},m^{*}_{q}\right)$ and $\left(l_{\mathrm{NGB}},M_{\mathrm{NGB}}\right)$ stands for the mass and orbital angular moment of constituent quarks and NGBs, respectively.
+The probability of each process to occur, allows us to compute the magnetic moment due to all possible transitions for a given valence quark.
+The orbital moments of the constituent quarks can also be written using the $\chi\mathrm{CQM}$ parameters $(a,\alpha,\beta,\zeta^{\prime})$, NGB masses $\left(M_{\pi},M_{\eta\prime},M_{K},M_{\eta}\right)$ and effective quark masses $\left(m^{*}_{u},m^{*}_{d},m^{*}_{s}\right)$, respectively as thirtyone ; fiftythree ; fiftyfour :
+${\left[\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\frac{3}{2}\frac{m_{u}^{*2}}{M_{\pi}\left(m_{u}^{*}+M_{\pi}\right)}-\frac{\alpha^{2}}{2}\frac{\left(M_{K}^{2}-3m_{u}^{*2}\right)}{M_{K}\left(M_{K}+m_{u}^{*}\right)}\right.+\frac{\beta^{2}}{6}\frac{M_{\eta}}{6\left(M_{\eta}+m_{u}^{*}\right)}$
+$\left.+\frac{\zeta^{\prime 2}}{3}\frac{M_{\eta^{\prime}}}{\left(m_{u}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{u},$
+${\left[\mu^{*}\left(d_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{3}{4}\frac{\left(M_{\pi}^{2}-2m_{d}^{*2}\right)}{M_{\pi}\left(m^{*}_{d}+M_{\pi}\right)}-\frac{\alpha^{2}}{2}\frac{M_{K}}{\left(m_{d}^{*}+M_{K}\right)}-\frac{\beta^{2}}{12}\frac{M_{\eta}}{\left(m_{d}^{*}+M_{\eta}\right)}\right.$ (39)
+$\left.-\frac{\zeta^{\prime 2}}{6}\frac{M_{\eta^{\prime}}}{\left(m_{d}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{d},$
+${\left[\mu^{*}\left(s_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{\alpha^{2}}{2}\frac{\left(M_{K}^{2}-3m_{s}^{*2}\right)}{M_{K}\left(m_{s}^{*}+M_{K}\right)}-\frac{\beta^{2}}{3}\frac{M_{\eta}}{\left(m_{s}^{*}+M_{\eta}\right)}-\frac{\zeta^{\prime 2}}{6}\frac{M_{\eta^{\prime}}}{\left(m_{s}^{*}+M_{\eta^{\prime}}\right)}\right]\mu_{s}^{*}.$
+In the present study, we have included the effect of relativistic quark confinement and have calculated the mass adjusted in-medium magnetic moments of its constituent quarks by the expression fiftyfive
+$\mu_{u}^{*}=2\left(1-\frac{\Delta M_{B}}{M_{B}^{*}}\right),\quad\mu_{d}^{*}=-\frac{1}{2}\mu_{u}^{*},\quad\mu_{s}^{*}=-\frac{m_{u}^{*}}{m_{s}^{*}}\left(1-\frac{\Delta M_{B}}{M_{B}^{*}}\right).$ (40)
+In the above relations, $\Delta M_{B}=M_{B,vac}-M_{B}^{*}$.
+Here, $M_{B,vac}$ and $M_{B}^{*}$ denotes the vacuum and in-medium mass of a given baryon, respectively.
+As discussed in the introduction of present manuscript, in Ref.
+thirtyone the impact of relativistic quark confinement was not considered and the simplified expression $\mu_{q}=\frac{e_{q}}{2m_{q}},q=u,d,s$ was used to calculate the magnetic moment of constituent quarks in the free space.
 To account for the contributions of orbital angular momentum polarizations ($\mu^{*L}_{B}$) towards the total baryonic magnetic moment, we have calculated the individual contributions from the valence quarks and quark sea as
@@
 In the current section we will also describe the expressions used in the calculation of sea quark spin polarization in a baryon, which are obtained by substituting the third component of the orbital angular momentum of sea quarks in place of each valence quark as,
-$q^{(\pm 1)}\rightarrow-T_{[q,GB]}q^{(\pm 1)}+\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}.\hskip 150.79968pt$ (49)
+$q^{(\pm 1)}\rightarrow-T_{[q,NGB]}q^{(\pm 1)}+\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}.\hskip 150.79968pt$ (42)
 In eq.
-49, $T_{[q,GB]}$ refers to the emission probability of a goldstone bosons (GB) from a quark $q^{(\pm 1)}$ and $\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}$ represents the transformation probability of a $q^{(\pm 1)}$ quark into two other quarks of $SU(3)$ model, given in thirtyone as
-$\left|\psi\left(u^{(\pm 1)}\right)\right|^{2}=ad^{(\pm 1)}+a\alpha^{2}s^{(\pm 1)},$ (50)
-$\left|\psi\left(d^{(\pm 1)}\right)\right|^{2}=au^{(\pm 1)}+a\alpha^{2}s^{(\pm 1)},$
-$\left|\psi\left(s^{(\pm 1)}\right)\right|^{2}=a\alpha^{2}u^{(\pm 1)}+a\alpha^{2}d^{(\pm 1)}.$
-The probability of transition a $q^{\pm 1}$ constituent quark into all the possible quarks ($u,d$ and $s$) along with GBs $\left(q\bar{q}^{\prime}\right)$ can be expressed as fiftyeight
-$P\left(u^{\pm}\rightarrow u\bar{q}^{\prime}+q^{\prime\mp}\right)=\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)u^{\mp}+ad^{\mp}+a\alpha^{2}s^{\mp},$ (51)
-$P\left(d^{\pm}\rightarrow d\bar{q}^{\prime}+q^{\prime\mp}\right)=au^{\mp}+\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)d^{\mp}+a\alpha^{2}s^{\mp},$
-$P\left(s^{\pm}\rightarrow s\bar{q}^{\prime}+q^{\prime\mp}\right)=a\alpha^{2}u^{\mp}+a\alpha^{2}d^{\mp}+\frac{a}{3}\left(2\beta^{2}+\zeta^{\prime 2}\right)s^{\mp}.$
-Using eq.
-51, we can easily obtain the $T_{[q,GB]}$ for each constituent quark.
-The valence and sea quark contributions of the orbital angular momentum part of the magnetic moment of the baryons, can thus be calculated easily.
-As an example, we have presented the orbital angular momentum contribution to baryonic magnetic moment of $\frac{1}{2}^{-}$ $n^{*}$ resonances as,
-$\mu^{*L}_{\mathrm{valence}}\left(n^{*}\right)=$ $\left(\frac{4}{9}\mu^{*}_{d}+\frac{2}{9}\mu^{*}_{u}\right)\cos^{2}\theta-\left(\frac{1}{9}\mu^{*}_{d}+\frac{2}{9}\mu^{*}_{u}\right)\sin^{2}\theta,$ (52)
-$\mu^{*L}_{\mathrm{sea}}\left(n^{*}\right)=$ $\frac{2a}{9}\left[-\left(1+2\alpha^{2}\right)\mu^{*}_{d}+\left(1-\alpha^{2}\right)\mu^{*}_{u}\right.\left.+3\alpha^{2}\mu^{*}_{s}\right]\cos^{2}\theta$
-$+\frac{a}{9}\left[-\left(1-\alpha^{2}\right)\mu^{*}_{d}\right.\left.+\left(1+2\alpha^{2}\right)\mu^{*}_{u}-3\alpha^{2}\mu^{*}_{s}\right]\sin^{2}\theta.$
-Using similar methodology as demonstrated above, the magnetic moments of the $\frac{1}{2}^{-}$ octet baryon resonances $p^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$ and $\Xi^{*0}$ as well as the low-lying $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$, and $S_{11}^{0}(1650)$ have been calculated within the $\chi\mathrm{CQM}$ framework.
-IV Results & Discussion
-We now discuss numerical results on the in-medium masses (section IV.1) and magnetic moments (section IV.2) of spin $\frac{1}{2}^{-}$ octet baryon resonances.
-The effective baryon masses have been listed in tables 1 and 2, so as to facilitate a discussion on the mass modifications.
-As an outcome of the calculations discussed in the previous sections, the vacuum and in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances have been illustrated in tables 3, 4, 5 and 6.
+42, $T_{[q,NGB]}$ refers to the emission probability of a Nambu-Goldstone bosons (NGB) from a quark $q^{(\pm 1)}$ and $\left|\psi\left(q^{(\pm 1)}\right)\right|^{2}$ represents the transformation probability of a $q^{(\pm 1)}$ quark into two other quarks of $SU(3)$ model, as given in Ref.
+thirtyone .
+Using the probability of transition of a constituent quark ($q^{\pm 1}$) into all other quarks ($u,d$ and $s$) accompanied by NGBs fiftyeight , we can easily obtain the $T_{[q,NGB]}$ for each constituent quark.
+The contributed values from the valence and sea quark polarizations of the orbital angular momentum part of the magnetic moment of the baryons, can thus be calculated easily.
+Working on the framework as described in this section, the magnetic moments of the $\frac{1}{2}^{-}$ octet baryon resonances $p^{*},n^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$ and $\Xi^{*0}$ as well as the low-lying $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$, and $S_{11}^{0}(1650)$ have been calculated within the $\chi\mathrm{CQM}$ framework.
+3 Results & Discussion
+We now discuss numerical results on the in-medium masses (section 3.1) and magnetic moments (section 3.2) of spin $\frac{1}{2}^{-}$ octet baryon resonances.
+Various parameters used in the current manuscript are listed in Table 1.
+The effective baryon masses have been listed in Tables 2 and 3, so as to facilitate a discussion on the mass modifications.
+As an outcome of the calculations discussed in the previous sections, the vacuum and in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances have been illustrated in Tables 4 to 7.
 For the sake of comparison with the baryonic magnetic moments obtained from earlier studies, we have tabulated the values calculated through different models in Table 8 for $\rho_{B}=0$.
-IV.1 In-medium masses of $\frac{1}{2}^{-}$ baryon resonances
-Figure 1: Effective masses of constituent quarks $u,d$ and $s$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-As the medium modifications of baryon masses is calculated using the medium modified masses of constituent quarks, we first discuss the impact of density and temperature on effective masses of quarks.
+As discussed in previous section, in our present work quark masses and energies are defined in terms of scalar and vector fields.
+The values of the confining parameter $k_{c}$ and couplings are fitted to binding energy value of $-16$ MeV at the nuclear saturation density.
+With this approach, at $\rho_{B}=0$, the constituent quark masses of $u$ and $d$ quarks are obtained to be 256 MeV while for s-quark it is 477 MeV.
+The difference in the constituent quark masses in present manuscript compared to Ref.
+thirtyone (for $u/d$ quarks 330 MeV and for $s$ quark 490 MeV), leads to different values for the magnetic moments of baryons in the free space as can be seen in Table 8.
+3.1 In-medium masses of $\frac{1}{2}^{-}$ baryon resonances
+Figure 1: In-medium masses of constituent quarks $u,d$ and $s$, at varying medium density $\rho_{B}$ (in $\rho_{0}$ units).
+As the in-medium baryon masses is determined using the medium modified masses of constituent quarks, we first discuss the impact of density and temperature on the effective quark masses.
 In fig.
-1, the in-medium masses, $m^{*}_{u},m^{*}_{d}$ and $m^{*}_{s}$, of the constituent quarks are shown as a function of density in units of nuclear saturation density, at $T=0$ MeV (on the left) and $T=100$ MeV (on the right)(where nuclear saturation density $\rho_{0}=0.16$ $\text{fm}^{-3}$).
-It also illustrates the in-medium modification in quark masses for different values of isospin asymmetry parameter ($i_{a}$) and strangeness fraction ($f_{s}$).
+1, the constituent quark masses, $m^{*}_{u},m^{*}_{d}$ and $m^{*}_{s}$, are shown to vary with the relative hadronic matter density $\rho_{B}/\rho_{0}$, at $T=0$ MeV (on the left) and $T=100$ MeV (on the right)(where nuclear saturation density $\rho_{0}=0.16$ $\text{fm}^{-3}$).
+It also illustrates the in-medium mass modifications of quarks, subject to a medium with finite isospin asymmetry ($i_{a}$) and strangeness fraction ($f_{s}$).
 From fig.
@@
 Thus an increase in the medium temperature causes a rise in the masses of constituent quarks, at the same density.
-The increase in isospin asymmetry of the medium leads to increase in-medium masses of constituent quarks.
+Keeping the medium density affixed, a rise in isospin asymmetry parameter leads to a surge in-medium masses of constituent quarks.
 A study on the basis of an extended NJL model has previously demonstrated that while the in-medium masses of light quarks approach their current quark masses, the $s$ quark condensate remains large even at densities close to $\rho_{B}\approx 1$ $\text{fm}^{-3}$ xia .
@@
 Using eq.
-25 the effective masses of $\frac{1}{2}^{-}$ baryon resonances are calculated in the isospin asymmetric strange matter, where the $E_{ispin}$ values are best fitted to procure the masses of $\frac{1}{2}^{-}$ baryon resonances in free space.
+26 the in-medium masses of $\frac{1}{2}^{-}$ baryon resonances are computed in the presence of a hadronic matter, where the $E_{ispin}$ values are best fitted to procure the masses of $\frac{1}{2}^{-}$ baryon resonances in free space.
 A quick glance at the arguments and figures depicts, a monotonous decrease in the in-medium baryon masses as a function of density.
@@
 As seen from fig.
-2, when symmetric nuclear matter density rises upto $\rho_{B}=3\rho_{0}$, the mass of $p^{*}$ resonance drops appreciably by 33.2%.
-At a density of $3\rho_{0}$, with finite isospin asymmetry $i_{a}=0.3$, the mass of $p^{*}$ is 1010.03 MeV which is a rise from 1003.60 MeV at $i_{a}=0$.
-In fact, with higher valence $d$ quark content in a baryon, the density effects are found to be suppressed.
-This can be seen from the 25.27% decrease in the effective mass of $n^{*}$, as the medium density reaches $\rho_{B}=3\rho_{0}$ from vacuum.
-Keeping the density unchanged, say at $\rho_{B}=3\rho_{0}$, the $M^{*}_{n}$ value is found to incur a marginal increase, as a function of asymmetry of the medium, to attain 1149.18(1147.02) MeV for $i_{a}=0.3(0)$.
+2, when symmetric nuclear matter density rises upto $\rho_{B}=3\rho_{0}$, the mass of $p^{*}$ resonance drops appreciably by a decrease of 25.27%.
+At a density of $3\rho_{0}$, with finite isospin asymmetry $i_{a}=0.3$, the mass of $p^{*}$ is 1149.18 MeV which is a mild rise from 1147.02 MeV at $i_{a}=0$.
+Taking into account a temperature rise from $T=0$ MeV to 100 MeV, we see distinctive baryonic mass curves for different strangeness fractions ($f_{s}=0,0.3$) of the medium.
+This exact same behaviour is seen to be replicated by the medium modified mass plots of $n^{*}$ as can be seen from fig.
+2, which is quite trivial as well considering the light quark compositions of $p^{*}$ and $n^{*}$.
 In contrary to light quark content baryons, for symmetric nuclear matter the mass modifications incurred by $\Sigma^{*}$ resonances, as seen in fig.
@@
 The extent of in-medium variations of the $N^{*}$ resonances is large compared to hyperons, like, in the case of $p^{*}$ and $n^{*}$ resonances.
-At a given value of $\rho_{B}$, with $i_{a}=0,T=0$ MeV, the impact of strangeness fraction, $f_{s}$, causes a sizable decrease in the masses of baryons.
+At a given value of $\rho_{B}$, with $i_{a}=0,T=0$ MeV, the influence of non-zero strangeness fraction, $f_{s}$, leads to a sizable decrease in the masses of baryons.
 Looking up the mass of $\Xi^{*-}$ from table 2, at $\rho_{B}=3\rho_{0},f_{s}=0.3(0)$, we observe that the value decreases to 1399.14(1439.75) MeV.
@@
 In the context of $\frac{1}{2}^{+}$ octet baryons, there has been studies on their in-medium mass modifications based on QMC hadron_mass , modified QMC (MQMC) mqmc , CQMF sixtytwo and also for decuplet, low-lying strange, charm, and bottom baryons in QMC framework tsushima1 ; tsushima2 .
-Figure 2: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-Figure 3: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-Figure 4: Effective masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-Figure 5: Effective masses of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-IV.2 In-medium magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances
-Figure 6: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-Figure 7: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-Figure 8: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-Figure 9: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-Figure 10: Effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-Figure 11: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-Figure 12: Effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-Figure 13: Contributions to effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark magnetic moments ($\mu_{orbit}^{*}$) are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density, $\rho_{0}$) at temperatures $T=0$ and 100 MeV.
-Results are shown for isospin asymmetry $i_{a}=0,0.3,0.5$ and strangeness fractions $f_{s}=0,0.3$.
-In this section we elaborately discuss the results of the vacuum and in-medium magnetic moments of ${\frac{1}{2}}^{-}$ octet baryon resonances $p^{*},n^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$, $\Xi^{*0}$ and low-lying $N^{*}$ resonances ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$, observed from figs.
+Figure 2: In-medium masses of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$, at varying medium density $\rho_{B}$ (in $\rho_{0}$ units).
+Figure 3: In-medium masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$, at varying medium density $\rho_{B}$ (in $\rho_{0}$ units).
+Figure 4: In-medium masses of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$, at varying medium density $\rho_{B}$ (in $\rho_{0}$ units).
+Figure 5: In-medium masses of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$, at varying medium density $\rho_{B}$ (in $\rho_{0}$ units).
+3.2 In-medium magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances
+Figure 6: Variation of effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units) with baryon density $\rho_{B}$ (in units of $\rho_{0}$).
+Figure 7: In-medium trends in explicit contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $p^{*}$ and $n^{*}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark ($\mu_{orbit}^{*}$) are presented.
+Figure 8: Variation of effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units) with baryon density $\rho_{B}$ (in units of $\rho_{0}$).
+Figure 9: In-medium trends in explicit contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Sigma^{*+}$, $\Sigma^{*-}$ and $\Sigma^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark ($\mu_{orbit}^{*}$) are presented.
+Figure 10: Variation of effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units) with baryon density $\rho_{B}$ (in units of $\rho_{0}$).
+Figure 11: In-medium trends in explicit contributions to effective magnetic moments of $\frac{1}{2}^{-}$ octet baryon resonances $\Xi^{*-}$ and $\Xi^{*0}$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark ($\mu_{orbit}^{*}$) are presented.
+Figure 12: Variation of effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units) with baryon density $\rho_{B}$ (in units of $\rho_{0}$).
+Figure 13: In-medium trends in explicit contributions to effective magnetic moments of $\frac{1}{2}^{-}$ $N^{*}$ resonances $S_{11}^{+}(1650)$ and $S_{11}^{0}(1650)$ (in $\mu_{N}$ units), from valance quarks ($\mu_{val}^{*}$), sea quarks ($\mu_{sea}^{*}$) and orbital moment of sea quark ($\mu_{orbit}^{*}$) are presented.
+This is where we elaborately examine the results of the vacuum and effective magnetic moments of ${\frac{1}{2}}^{-}$ octet baryon resonances $p^{*},n^{*},\Sigma^{*+},\Sigma^{*-},\Sigma^{*0},\Xi^{*-}$, $\Xi^{*0}$ and low-lying $N^{*}$ resonances ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$, observed from figs.
 6, 7, 8, 9, 10, 11, 12 and 13.
@@
 The observed in-medium behavior of $\mu^{*}_{p^{*}}$ and $\mu^{*}_{n^{*}}$, shows a steeper slope at lower densities of nuclear matter ($\rho_{B}\leq 3\rho_{0}$) and gradually attains saturation at much higher densities.
-As we trace the plot of these effective baryon magnetic moments while the medium density rises to $\rho_{B}=3\rho_{0}$, we observe that $\mu^{*}_{p^{*}}$ increases by $0.59\mu_{N}$ and $\mu^{*}_{n^{*}}$ initially approaches a minima at $-1.382\mu_{N}$ (at $\rho_{B}\approx\rho_{0}$) and then climbs up close to its vacuum value at $-1.375\mu_{N}$.
+As we trace the plot of these effective baryon magnetic moments (at $T=0$) while the symmetric medium density rises to $\rho_{B}=3\rho_{0}$, we observe that $\mu^{*}_{p^{*}}$ increases by $0.19\mu_{N}$ and $\mu^{*}_{n^{*}}$ initially approaches a minima at $-1.382\mu_{N}$ (at $\rho_{B}\approx 1.1\rho_{0}$) and then climbs up close to its vacuum value at $-1.376\mu_{N}$.
 A conclusive understanding of the observed effective baryonic magnetic moments of $p^{*}$ and $n^{*}$ can be provided by analyzing fig.
 7, where the contributions from valence quarks, sea quarks and orbital angular momentum of the quark sea have been shown individually.
-For example, in case of $n^{*}$ in symmetric nuclear matter, the decrease in the valence quark contributions dominates over the increasing contributions of sea quarks and the orbital contributions of sea quarks, which leads to an overall decreasing behavior of $\mu^{*}_{n^{*}}$ below the saturation density of the medium ($\rho_{0}$).
-As the medium density moves past this point ($\rho_{B}\approx\rho_{0}$), owing to the dominance of $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$ contributions the effective in-medium behavior of $\mu^{*}_{n^{*}}$ reverses and it approaches higher values.
+For instance, take $n^{*}$ in symmetric nuclear matter, the decrease in the valence quark contributions dominates over the increasing contributions of sea quarks and the orbital contributions caused due to them, which leads to an overall decreasing behavior of $\mu^{*}_{n^{*}}$ below the saturation density of the medium ($\rho_{0}$).
+As the medium density moves past this point ($\rho_{B}\approx\rho_{0}$), owing to the dominance of $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$ contributions, the effective behavior of $\mu^{*}_{n^{*}}$ reverses and it approaches higher values.
 On a similar note, the density dependency of $\mu^{*}_{p^{*}}$ can be comprehended by taking into account the increase in $\mu^{*}_{val}$ dominating over the decrease in $\mu^{*}_{sea}$ and $\mu^{*}_{orbit}$.
@@
 7.
-As we have included “Goldstone Boson Exchange”(GBE) effect in this study, it is important to note that the sea quark effect has an opposite sign as compared to the valence quark effect, whereas the orbital angular momentum of sea quarks has the same contribution as that from the valence quark effect (see Table 3).
+As we have included “Nambu Goldstone Boson Exchange”(NGBE) effect in this study, it is critical to note that the influence of sea quark leads to an opposite sign as compared to the valence quark effect, while on the other hand the orbital angular momentum of sea quarks has the similar contribution as that caused by the valence quarks (see Table 4).
 The sign of these contributions can be justified, if we realize that the production of the quark sea is associated with a spin flip process (see eq.
-30) and hence, the sea quarks has to have an opposite effect.
-An earlier study highlights the impact of Goldstone Boson (GB) degrees in chiral quark model ($\chi$QM), where the magnetic moments calculated for $p^{*}$ and $n^{*}$ in vacuum, by considering GBE interaction, are 1.4 $\mu_{N}$ and $-0.9\mu_{N}$ for $\chi$QM, which are slightly lower than the corresponding values, 1.7 $\mu_{N}$ and $-1.1\mu_{N}$ as obtained from CQM framework liu .
-In addition, the magnetic moments procured from $\chi$QM by considering Goldstone Boson Exchange (GBE) and One Gluon Exchange (OGE), separately, gives similar magnetic moment values and are in close agreement to our vacuum state results liu .
+32) and hence, the sea quarks has to have an opposite effect.
+An earlier study highlights the impact of Nambu-Goldstone Boson (NGB) degrees in chiral quark model ($\chi$QM), where the magnetic moments calculated for $p^{*}$ and $n^{*}$ in vacuum, by considering NGBE interaction, are 1.4 $\mu_{N}$ and $-0.9\mu_{N}$ for $\chi$QM, which are slightly lower than the corresponding values, 1.7 $\mu_{N}$ and $-1.1\mu_{N}$ as obtained from CQM framework liu .
+In addition, the magnetic moments procured from $\chi$QM by considering Nambu Goldstone Boson Exchange (NGBE) and One Gluon Exchange (OGE), separately, gives similar magnetic moment values and are in close agreement to our vacuum state results liu .
 Two other studies in the literature, with quasibound state approach quasibound and chiral unitary model chiral_unitary have previously calculated for magnetic moments of these two resonances and could be compared from table 8.
-Table 7 facilitates a comparitive study of the magnetic moments of $p^{*}$ and $n^{*}$ resonances obtained from various models.
+Table 8 facilitates a comparative analysis of the magnetic moments of $p^{*}$ and $n^{*}$ resonances obtained from various models.
 Figure 8 depicts the in-medium modifications of $\mu^{*}_{\Sigma^{*+}}$, $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$, as a function of $\rho_{B}/\rho_{0}$ and the behavior of the individual contributions have been plotted in fig.
@@
 From a point of comparison if you see Table 8, the magnetic moments of $\mu^{*}_{\Sigma^{*+}}$ (1.647 $\mu_{N}$), $\mu^{*}_{\Sigma^{*-}}$ $(-0.754\mu_{N})$ and $\mu^{*}_{\Sigma^{*0}}$ (0.446 $\mu_{N}$) in free space are found to be in agreement with the NCQM results 1.814 $\mu_{N}$, $-0.689\mu_{N}$ and 0.820 $\mu_{N}$, respectively.
-As seen from Tables 3 and 5, for $f_{s}=0(0.3)$ at $\rho_{B}=3\rho_{0}$, $\mu^{*}_{\Sigma^{*+}}$ and $\mu^{*}_{\Sigma^{*-}}$ values attain 1.697(1.723) $\mu_{N}$ and $-0.770(-0.779)\mu_{N}$.
+As seen from Tables 4 and 6, for $f_{s}=0(0.3)$ at $\rho_{B}=3\rho_{0}$, $\mu^{*}_{\Sigma^{*+}}$ and $\mu^{*}_{\Sigma^{*-}}$ values attain 1.697(1.723) $\mu_{N}$ and $-0.770(-0.779)\mu_{N}$.
 This suggests that as we introduce finite strangeness fraction to the non-strange medium, there will be an enhancement in the nature of in-medium variations of baryonic magnetic moments of $\Sigma^{*+}$, $\Sigma^{*0}$ and $\Sigma^{*-}$.
 For finite $f_{s}$, the magnetic moments of $\Sigma^{*+}$ and $\Sigma^{*0}$ surges while that of $\Sigma^{*-}$ approaches lower values, instead of attaining a saturation at large densities.
-For an instance, the in-medium values of $\mu^{*}_{\Sigma^{*+}}$, $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$ (in units of $\mu_{N}$) at $i_{a}=0(0.5)$, when referred to Tables 3 and 4, are $1.697(1.698),-0.770(-0.763)and0.463(0.472)$ in a medium density of $3\rho_{0}$.
+For an instance, the in-medium values of $\mu^{*}_{\Sigma^{*+}}$, $\mu^{*}_{\Sigma^{*-}}$ and $\mu^{*}_{\Sigma^{*0}}$ (in units of $\mu_{N}$) at $i_{a}=0(0.5)$, when referred to Tables 4 and 5, are $1.697(1.698),-0.770(-0.763)and0.463(0.472)$ in a medium density of $3\rho_{0}$.
 This and a glance into fig.
@@
 Such $i_{a}$ dependent fine shifts in the baryonic magnetic moment have been previously found in works, in the present formalism, dedicated to octet baryons fiftythree and decuplet baryons fiftyfour .
-We find it interesting to note that for $\Xi^{*}$ baryons the density effects on $\mu^{*}_{\Xi^{*-}}$ and $\mu^{*}_{\Xi^{*0}}$ values show impressive similarity, owing to the valence quark spin polarizations and corresponding orbital moments of two $s$ quarks that suppresses the impact of light $u$ and $d$ quarks in $\Xi^{*-}$ and $\Xi^{*0}$, respectively sixtytwo .
+We find it interesting to note that for $\Xi^{*}$ baryons the density effects on $\mu^{*}_{\Xi^{*-}}$ and $\mu^{*}_{\Xi^{*0}}$ values show impressive similarity, owing to the valence quark spin polarizations and corresponding orbital moments of two $s$ quarks that suppresses the impact of light quarks present in them sixtytwo .
 The baryonic magnetic moment of $\Xi^{*-}$, as seen from fig.
-10, rises steeply with increase in the relative density ($\rho_{B}/\rho_{0}$) of the medium.
+10, rises steeply with increasing relative hadronic matter density ($\rho_{B}/\rho_{0}$).
 Figure 11 clearly illustrates that for $\Xi^{*-}$, the cumulative increasing behavior of $\mu^{*}_{val}$ and $\mu^{*}_{sea}$ contributions overshadows the decreasing orbital contributions from sea quarks ($\mu^{*}_{orbit}$).
@@
 From fig.
-12, we observe that the magnetic moments of ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$ increases, overall as a function of density of the hadronic matter $\rho_{B}$ (in $\rho_{0}$ units).
+12, we observe that the magnetic moments of ${S^{+}_{11}}(1650)$ and ${S^{0}_{11}}(1650)$ increases, overall as a function of relative medium density $\rho_{B}/\rho_{0}$.
 As from fig.
@@
 seventyfour have discussed a strong correlation of the baryonic magnetic moment with that of the bag radius of a baryon, as the two quantities have a similar behavior as density dependent functions.
-Since this is the first work on medium modifications of $\frac{1}{2}^{-}$ baryon resonances, so referring to previous works on in-medium density effects on magnetic moments of baryons using the current formalism sixtytwo ; fiftyfour and QMC tsushima2 , might facilitate the purpose of a comparative study for future work on the magnetic moments of $\frac{1}{2}^{-}$ baryon resonances.
-V Summary
-To summarize, in the present work we have investigated the modifications induced on the masses as well as the magnetic moments of $J_{p}={\frac{1}{2}}^{-}$ baryon resonances in isospin asymmetric strange hadronic matter.
-In light of chiral $SU(3)$ quark mean field model the in-medium effects are introduced and we computed the effective baryon masses by using the masses of constituent quarks.
+Since this is the first work on medium modifications of $\frac{1}{2}^{-}$ baryon resonances, so referring to previous works on in-medium density effects on magnetic moments of baryons using the current formalism sixtytwo ; fiftyfour and QMC tsushima2 , might facilitate the purpose of a comparative study for future work on the baryonic magnetic moments of $\frac{1}{2}^{-}$ octet resonances.
+The effective masses of positive parity octet baryons are also observed to decrease as a function of nuclear density.
+As an example, the effective masses of proton $p$ is observed to decrease by 44 % as baryon density is decreased from 0 to $3\rho_{0}$ and this can be compared with the decrease of 25.27% for negative parity $p^{*}$ baryon.
+At high baryon density chiral symmetry is expected to restore partially and in this limit, properties of positive and negative parity baryons may converge to same values.
+This kind of picture can be seen more clearly within the present model calculations, if negative parity baryons are considered along with the positive parity baryons in the solution of system of non-linear equations corresponding to scalar and vector densities of baryons.
+As we discussed earlier, in the present work properties of negative parity baryons are investigated in strange matter comprising of positive parity nucleons and hyperons.
+In our future work, we shall extend our present model calculations to study the thermodynamics of dense matter considering explicitly both positive and negative baryons in equations of motion of scalar and vector fields through the definitions of scalar and vector densities of baryons.
+4 Summary
+In a nutshell, this work investigates the medium modified trends induced on the masses and moreover on the magnetic moments of $J_{p}={\frac{1}{2}}^{-}$ baryon resonances in isospin asymmetric strange hadronic matter.
+The effects caused due to the presence of a medium are introduced through the CQMF model, following which we computed the effective baryon masses by using the masses of constituent quarks.
 The mass adjusted quark magnetic moments are then calculated, to include relativistic quark confinement, by using the baryon masses as inputs.
 In hadronic medium the in-medium masses of lights quarks, $u$ $d$, vary significantly in comparison to the observed changes in medium modified masses of $s$ quark.
-In the $\chi$CQM formalism, we then computed the baryonic magnetic moments from the contributions due to valence quarks, sea quarks and orbital angular momentum due of sea quarks.
+In the $\chi$CQM formalism, we then obtained the baryonic magnetic moments by the summation over the contributed values coming from valence quarks, sea quarks and orbital angular momentum due of sea quarks.
 As derived from the results is that, the medium modifications are more prominent for the baryons comprised of light $u$ or $d$ quarks, than for hyperons.
-For strange matter, the in-medium constituent quark masses, effective baryonic masses as well as the effective baryonic magnetic moments suffer appreciable enhancement in their in-medium behaviour.
+For strange matter, the in-medium masses of constituent quark and baryons as well as the effective baryonic magnetic moments suffer appreciable enhancement in their in-medium behaviour.
 By carrying out this study at finite temperatures and finite $i_{a}$ values of the hadronic medium, we have noted a fine shift in the in-medium baryonic magnetic moments, although the temperature effects have been negligible below $T=150$ MeV.
+The results from this study could be a point of reference to experiments which study hadron spectroscopy by probing photoproductions.
+In Jefferson Lab the GlueX experiment has measured the width of the N(1535) state to be narrower (less than 100 MeV) making it undetetctable, whereas its cascade program has aligned its goals to explore the negative parity excited $\Xi$ baryonsbaryon_spectroscopy .
+In the limelight of excited cascade resonances, a report from the Belle Collaboration sheds light on the detection of two states in low mass regime of which one ought to be $\Xi$(1690) Belle .
+Such experimental studies have research interests directed towards a detailed understanding of QCD for low-lying hadronic structures.
+Hence, the density dependent study of negative parity first excited states of octet baryon resonances is a crucial point of experimental interest at exploring the predictions of constituent quark models.
+Additionally, this work has scope for further extrapolation in the context of higher density astrophysical objects like neutron stars, to understand the impact of a dense medium on the baryonic magnetic moments medium_effect_neutron .
+$\begin{array}[]{|c|c|c|c|c|c|c|}\hline\cr k_{0}&k_{1}&k_{2}&k_{3}&k_{4}\\ \hline\cr 4.94&2.12&-10.16&-5.38&-0.06\\ \hline\cr\sigma_{0}\,(\text{MeV})&\quad\zeta_{0}\,(\text{MeV})&\quad\chi_{0}\,(\text{MeV})&\quad\xi&\quad\rho_{0}\,(\text{fm}^{-3})\\ \hline\cr-92.8&-96.5&254.6&\frac{6}{33}&0.16\\ \hline\cr g_{\sigma}^{u}=g_{\sigma}^{d}&\quad g_{\sigma}^{s}=g_{\zeta}^{u}=g_{\zeta}^{d}&\quad g_{\delta}^{u}&g_{\zeta}^{s}=g_{s}&g_{4}\\ \hline\cr 2.72&0&2.72&3.847&37.4\\ \hline\cr g_{\rho}^{p}=g_{\rho}^{u}&g_{\omega}^{N}=3g_{\omega}^{u}&\quad g_{\rho}^{p}&\quad m_{\pi}\,(\text{MeV})&\quad m_{K}\,(\text{MeV})\\ \hline\cr 2.72&9.69&8.886&139&494\\ \hline\cr\end{array}$
+Table 1: In above table various parameters used in the current manuscript are summarized.
 Baryon Vacuum Mass $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
@@
 $M_{S^{0}_{11}(1650)}^{*}$ 1650.0 $1439.01$ $1439.98$ $1435.09$ $1438.69$ $1263.40$ $1265.55$ $1254.90$ $1257.48$
-$M_{p^{*}}^{*}$ 1535.0 $1241.14$ $1245.05$ $1225.47$ $1233.20$ $1003.60$ $1010.03$ $953.84$ $962.62$
+$M_{p^{*}}^{*}$ 1535.0 $1323.31$ $1326.33$ $1319.37$ $1325.08$ $1147.02$ $1151.91$ $1138.50$ $1143.92$
 $M_{n^{*}}^{*}$ 1535.0 $1323.31$ $1324.28$ $1319.37$ $1322.98$ $1147.02$ $1149.18$ $1138.50$ $1141.08$
@@
 $M_{S^{0}_{11}(1650)}^{*}$ 1650.0 $1469.66$ $1469.70$ $1454.68$ $1453.87$ $1289.10$ $1289.73$ $1270.06$ $1269.38$
-$M_{p^{*}}^{*}$ 1535.0 $1283.87$ $1286.17$ $1251.81$ $1253.30$ $1040.65$ $1045.23$ $972.85$ $976.01$
+$M_{p^{*}}^{*}$ 1535.0 $1354.07$ $1355.90$ $1339.04$ $1340.30$ $1172.82$ $1176.51$ $1153.71$ $1156.32$
 $M_{n^{*}}^{*}$ 1535.0 $1354.07$ $1354.10$ $1339.04$ $1338.22$ $1172.82$ $1173.45$ $1153.71$ $1153.02$
@@
 ${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.065 -0.286 0.125 0.904 1.213 -0.334 0.064 0.943
-${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.299 -0.392 0.207 2.114 2.843 -0.483 0.016 2.376
+${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.156 -0.368 0.194 1.982 2.487 -0.423 0.014 2.078
 ${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.464 0.263 -0.181 -1.382 -1.689 0.305 0.007 -1.376
@@
 ${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.833 -0.282 -0.084 -1.200 -0.687 -0.318 -0.019 -1.026
-Table 3: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric nuclear matter ($i_{a}=0$ and $f_{s}=0$) at a medium temperature $T=0$ MeV.
+Table 4: Computed values of effective magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric nuclear matter ($i_{a}=0$ $f_{s}=0$) at $T=0$ MeV.
 Baryons $\rho_{B}=0$ $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
@@
 ${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.062 -0.285 0.134 0.911 1.205 -0.331 0.079 0.953
-${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.280 -0.389 0.213 2.104 2.792 -0.475 0.032 2.349
-${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.459 0.263 -0.185 -1.382 -1.676 0.303 -0.003 -1.376
+${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.151 -0.367 0.196 1.980 2.477 -0.421 0.019 2.074
+${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.463 0.263 -0.183 -1.382 -1.686 0.305 0.004 -1.376
 ${\Sigma}^{*+}\left(\mu_{N}\right)$ 1.761 -0.343 0.229 1.647 1.927 -0.380 0.137 1.684 2.072 -0.416 0.042 1.698
@@
 ${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.840 -0.281 -0.087 -1.209 -0.710 -0.314 -0.026 -1.051
-Table 4: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric nuclear matter ($i_{a}=0.5$ and $f_{s}=0$) at a medium temperature $T=0$ MeV.
+Table 5: Computed values of effective magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric nuclear matter ($i_{a}=0.5$ $f_{s}=0$) at $T=0$ MeV.
 Baryons $\rho_{B}=0$ $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
@@
 ${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.068 -0.287 0.124 0.904 1.221 -0.337 0.062 0.946
-${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.328 -0.397 0.205 2.136 2.991 -0.509 0.005 2.487
+${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.162 -0.369 0.191 1.984 2.506 -0.426 0.004 2.084
 ${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.468 0.264 -0.178 -1.382 -1.701 0.308 0.018 -1.375
@@
 ${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.841 -0.285 -0.083 -1.209 -0.703 -0.328 -0.016 -1.048
-Table 5: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric hyperonic matter ($i_{a}=0$ and $f_{s}=0.3$) at a medium temperature $T=0$ MeV.
+Table 6: Computed values of effective magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of symmetric hyperonic matter ($i_{a}=0$ $f_{s}=0.3$) at $T=0$ MeV.
 Baryons $\rho_{B}=0$ $\rho_{B}=\rho_{0}$ $\rho_{B}=3\rho_{0}$
@@
 ${S^{0}_{11}(1650)}\left(\mu_{N}\right)$ 0.929 -0.244 0.191 0.875 1.061 -0.284 0.135 0.911 1.216 -0.334 0.075 0.957
-${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.295 -0.392 0.216 2.119 2.944 -0.501 0.016 2.459
+${p}^{*}\left(\mu_{N}\right)$ 1.859 -0.318 0.349 1.889 2.142 -0.365 0.201 1.977 2.485 -0.423 0.013 2.075
 ${n}^{*}\left(\mu_{N}\right)$1.262 0.226 -0.339 -1.375 -1.458 0.262 -0.187 -1.382 -1.694 0.306 0.012 -1.375
@@
 ${\Xi}^{*0}\left(\mu_{N}\right)$0.903 -0.248 -0.149 -1.301 -0.851 -0.282 -0.088 -1.222 -0.722 -0.324 -0.022 -1.069
-Table 6: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric hyperonic matter ($i_{a}=0.5$ and $f_{s}=0.3$) at a medium temperature $T=0$ MeV.
+Table 7: Computed values of in-medium magnetic moments of $\frac{1}{2}^{-}$ baryon resonances at different medium densities ($\rho_{B}/\rho_{0}=0,1,3$) are tabulated in the presence of asymmetric hyperonic matter ($i_{a}=0.5$ $f_{s}=0.3$) at $T=0$ MeV.
 Baryons $\chi$CQM $\chi$QMliu CQMliu NCQMthirtyone Quasiboundquasibound CUchiral_unitary
-$\mu_{\text{val}}$ $\mu_{\text{sea}}$ $\mu_{\text{orbit}}$ $\mu_{B}$ $\mu_{B}$ (GBE) $\mu_{B}$ (OGE) $\mu_{B}$ (GBE) $\mu_{B}$ $\mu_{B}$ $\mu_{B}$
+$\mu_{\text{val}}$ $\mu_{\text{sea}}$ $\mu_{\text{orbit}}$ $\mu_{B}$ $\mu_{B}$ (NGBE) $\mu_{B}$ (OGE) $\mu_{B}$ (NGBE) $\mu_{B}$ $\mu_{B}$ $\mu_{B}$
 ${S^{+}_{11}(1650)}\left(\mu_{N}\right)$ 0.140 -0.169 -0.136 -0.165 0.1 0.0 0.3 0.106 - -
@@
 Table 8: The above table allows a comparative study of magnetic moments of $\frac{1}{2}^{-}$ baryon resonances, computed using different models in the free space ($\rho_{B}=0$).
-VI APPENDIX
-By minimizing the thermopotential in eq.
-27, we derive the equations of motion for scalar fields $\sigma$, $\zeta$, the dilaton field $\chi$, scalar iso-vector field $\delta$ and the vector fields $\omega,\rho,\phi$ in view of eq.
-29 as in fiftyfour
-[…]
-$\,+\,\frac{\chi^{2}}{\chi_{0}^{2}}m_{\pi}^{2}f_{\pi}-\left(\frac{\chi}{\chi_{0}}\right)^{2}m_{\omega}\omega^{2}\frac{\partial m_{\omega}}{\partial\sigma}\,-\left(\frac{\chi}{\chi_{0}}\right)^{2}m_{\rho}\rho^{2}\frac{\partial m_{\rho}}{\partial\sigma}\,+\,\sum_{i}\frac{\partial M_{i}^{\ast}}{\partial\sigma}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}$ (53)
-[…]
-$+\,\frac{\chi^{2}}{\chi_{0}^{2}}\left(\sqrt{2}m_{K}^{2}f_{K}\,-\,\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\right)+\,\sum_{i=\Lambda,\Sigma^{\pm,0},\Xi^{-,0}}\frac{\partial M_{i}^{\ast}}{\partial\zeta}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}=0,$ (54)
-$\frac{\partial\Omega}{\partial\chi}=k_{0}\chi\left(\sigma^{2}+\zeta^{2}+\delta^{2}\right)-k_{3}\left(\sigma^{2}-\delta^{2}\right)\zeta+\frac{2\chi}{\chi_{0}^{2}}\left[m_{\pi}^{2}f_{\pi}\sigma+\left(\sqrt{2}m_{K}^{2}f_{K}\,-\,\frac{1}{\sqrt{2}}m_{\pi}^{2}f_{\pi}\right)\zeta\right]$
-$-\frac{\chi}{\chi_{0}^{2}}\left(m_{\omega}^{2}\omega^{2}+m_{\rho}^{2}\rho^{2}\right)+\left(4k_{4}+1+{\rm ln}\frac{\chi^{4}}{\chi_{0}^{4}}-\frac{4\xi}{3}{\rm ln}\left(\left(\frac{\left(\sigma^{2}-\delta^{2}\right)\zeta}{\sigma_{0}^{2}\zeta_{0}}\right)\right)\right)\chi^{3}=0,$ (55)
-[…]
-$-\,\sum_{i}g_{\delta i}\rho_{i}^{s}=0,~{}~{}~{}~{}~{}$ (56)
-$\frac{\partial\Omega}{\partial\omega}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\omega}^{2}\omega+4g_{4}\omega^{3}+12g_{4}\omega\rho^{2}-\sum_{i}g_{\omega i}\rho_{i}=0,$ (57)
-$\frac{\partial\Omega}{\partial\rho}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\rho}^{2}\rho+4g_{4}\rho^{3}+12g_{4}\omega^{2}\rho-\sum_{i}{g_{\rho i}}\rho_{i}=0,$ (58)
-and
-$\frac{\partial\Omega}{\partial\phi}=\left(\frac{\chi^{2}}{\chi_{0}^{2}}\right)m_{\phi}^{2}\phi+8g_{4}\phi^{3}-\sum_{i}g_{\phi\,i}\rho_{i}=0.$ (59)
-In eqs. 57, 58 and 59 $\rho_{i}$ represents the vector density of baryons and is given by,
-$\rho_{i}=\gamma_{i}\int\frac{d^{3}k}{(2\pi)^{3}}\Big{(}f_{i}(k)-\bar{f}_{i}(k)\Big{)},$ (60)
-and in sections VI, VI and VI $\rho_{i}^{s}$ stands for the scalar density of baryons, given as
-$\rho_{i}^{s}=\gamma_{i}\int\frac{d^{3}k}{(2\pi)^{3}}\frac{M_{i}^{*}}{E^{\ast}_{i}(k)}\Big{(}f_{i}(k)+\bar{f}_{i}(k)\Big{)}.$ (61)
-Here, $f_{i}(k)$ and $\bar{f}_{i}(k)$ appearing in eqs. 60 and 61 represent the Fermi distribution functions for fermions and anti-fermions, at finite temperature and are expressed as
-$f_{i}(k)$ $=$ $\frac{1}{1+\exp\left[(E^{*}_{i}(k)-\nu^{*}_{i})/k_{B}T\right]}~{}~{}~{}\text{and}~{}~{}~{}\bar{f}_{i}(k)=\frac{1}{1+\exp\left[(E^{*}_{i}(k)+\nu^{*}_{i})/k_{B}T\right]}~{}.$ (62)